### v1

# Linearization of ergodic McKean SDEs and applications

Abstract
In this article, we consider McKean stochastic differential equations, as well as their corresponding McKean–Vlasov partial differential equations, which admit a unique stationary state, and we study the linearized Itô diffusion process that is obtained by replacing the law of the process in the convolution term with the unique invariant measure. We show that the law of this linearized process converges to the law of the nonlinear process exponentially fast in time, both in relative entropy and in Wasserstein distance. We study the problem in both the whole space and the torus. We then show how we can employ the resulting linear (in the sense of McKean) Markov process to analyze properties of the original nonlinear and nonlocal dynamics that depend on their long-time behavior. In particular, we propose a linearized maximum likelihood estimator for the nonlinear process which is asymptotically unbiased, and we study the joint diffusive-mean field limit of the underlying interacting particle system.
AMS subject classifications. 26D10, 35Q70, 35Q83, 60J60, 82C22.
Key words. Diffusive limit, interacting particle systems, invariant measure, logarithmic Sobolev inequality, maximum likelihood estimation, McKean SDE, McKean–Vlasov PDE, relative entropy.
1 Introduction
Interacting particle systems and their mean field limit have been an active area of research in the last few decades, due to their many applications to problems in physics [6, 31], biology [58], social sciences [19, 28, 30], algorithms for sampling and optimization [15, 16] and, more recently, machine learning [57, 42, 54, 12]. Starting from earlier fundamental works, e.g., [46, 59], in recent years there have been important improvements on the understanding of uniform propagation of chaos and its connection to the logarithmic Sobolev inequality (LSI) [41, 13, 39, 33, 43, 21, 23], nonuniqueness of stationary states [44], and phase transitions [14, 22], as well as the study of fluctuations around the mean field limit [37] and the Dean–Kawasaki equation [20, 35, 38, 29]. Moreover, there have been several studies on the inference problem for the McKean stochastic differential equation (SDE) and the McKean–Vlasov partial differential equation (PDE). Different approaches, based on maximum likelihood [24], stochastic gradient descent in continuous time [56], spectral theoretic (e.g., martingale eigenfunction estimator) techniques [51], deconvolution techniques [1], the generalized method of moments [52, 18], have been introduced and analyzed. In addition, fully Bayesian nonparametric inverse problems for the McKean–Vlasov PDE have been considered in [45]. An interesting aspect of the inference problem for the McKean SDE, when the measurement model consists of observations of the particle system, is the nonuniqueness of stationary states for the mean field PDE at high interaction strengths (equivalently, low temperatures) if the confining and interaction potentials are not convex on the whole space [62] or not $H$-stable on the torus [14, 22]. This makes the use of inference methodologies that rely on the ergodic properties of the dynamics problematic. This issue was addressed in [51] for the Desai–Zwanzig model by first inferring the correct invariant measure and then linearizing the generator of the McKean dynamics around it, and in [45] by considering observations only at short times, for appropriately chosen initial conditions. One of the main goals of this paper is to further develop the linearization methodology introduced in [51], in the context of parametric inference for McKean SDEs.
In addition, multiscale problems for interacting particle systems and McKean SDEs have been studied in recent years [32, 22, 4]. In particular, in [22] the joint diffusive-mean field limit for systems of weakly interacting reversible diffusions with periodic confining and interaction potentials was studied. It was shown that the diffusive and mean field limits commute when the mean field system restricted on the torus has a unique invariant measure. On the other hand, it was rigorously established that the two limits do not commute when the dynamics on the torus exhibits phase transitions. Furthermore, the covariance matrix of the limiting Brownian motion can be calculated using the standard homogenization (Kipnis–Varadhan) formula [49, Section 13.6.1], for a linear elliptic cell problem, obtained by calculating the convolution in the generator of the McKean SDE with respect to the Gibbs measure of the dynamics in the torus. In the present work, we show that the the diffusive limit of the interacting particle system and the formula for the diffusion coefficient can be obtained from the linearization procedure.
In this paper we consider the dynamics of a system of $N$ weakly interacting reversible Langevin diffusions, described by the following system of Itô SDEs
$\mathrm{d}X_{t}^{(n)}$ $=-\nabla V(X_{t}^{(n)})\,\mathrm{d}t-\frac{1}{N}\sum_{i=1}^{N}\nabla W(X_{t}^{(n)}-X_{t}^{(i)})\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t}^{(n)},$ (1.1)
$X_{0}^{(n)}$ $\sim\mu_{0},\qquad n=1,\dots,N,$
where $V,W\in\mathcal{C}^{2}(\mathbb{D})$ are the confining and interacting potentials, respectively, $\beta>0$ is the inverse temperature, and $\mu_{0}\in\mathcal{P}(\mathbb{D})$ is the initial distribution which is common among all the particles. The domain $\mathbb{D}$ can be the entire space $\mathbb{R}^{d}$ or the multidimensional torus $\mathbb{T}^{d}$. Under suitable conditions on the confining and interaction potentials, a rather complete description of the propagation of chaos phenomenon has been achieved, as well as of the properties of the mean field PDE [41, 39, 23]. In particular, in the limit as $N\to\infty$, particles become decorrelated and $X_{t}^{(n)}$, $n=1,\dots,N$, converge to the mean field limit
$\mathrm{d}X_{t}=-\nabla V(X_{t})\,\mathrm{d}t-(\nabla W*f_{t})(X_{t})\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},\qquad X_{0}\sim\mu_{0},$ (1.2)
where $f_{t}\colon\mathbb{D}\to\mathbb{R}$ denotes the density of the law $\mu_{t}$ of the process $X_{t}$ with respect to the Lebesgue measure, which satisfies the McKean–Vlasov PDE
$\frac{\partial f_{t}}{\partial t}=\nabla\cdot(\nabla Vf_{t}+(\nabla W*f_{t})f_{t})+\beta^{-1}\Delta f_{t}.$ (1.3)
The, possibly more than one, invariant measures $\mathrm{d}\mu_{\infty}(x)=f_{\infty}(x)\,\mathrm{d}x$ of the McKean SDE (1.2) are solutions of the Kirkwood–Monroe integral equation [60, 25]
$f_{\infty}=\frac{1}{Z}e^{-\beta(V+(W*f_{\infty}))},\qquad Z=\int_{\mathbb{D}}e^{-\beta(V(z)+(W*f_{\infty})(z))}\,\mathrm{d}z.$ (1.4)
It is well known that, under the conditions on the interaction and confining potentials that ensure uniform propagation of chaos for the interacting particle system, the Mckean SDE is geometrically ergodic, i.e., it converges exponentially fast in $L^{1}$, relative entropy, and 2–Wasserstein distance to its unique invariant measure (see, e.g., [41] for the dynamics in the whole space $\mathbb{R}^{d}$). Moreover, a different approach is presented in [27], where the authors prove convergence to equilibrium in Wasserstein distance by means of coupling techniques, specifically combination of reflection and synchronous couplings, which enable them to obtain explicit contraction rates. Recent developments in the study of uniform propagation of chaos and of its connection to the uniform LSI can be found in [39, 23]. This idea has also been used in [33] to prove the propagation of chaos on the torus under different assumptions on the interaction potential that replace convexity. Considering the multidimensional torus $\mathbb{T}^{d}$ as a state space, a different approach based on the (Lions) derivatives of the semigroup generated by the Fokker–Planck equation is proposed in [21].
Let us now place ourselves in the setting where we have uniform propagation of chaos and, therefore, the mean field dynamics converges exponentially fast to the unique invariant measure. In this case, the McKean SDE (1.2) and the linear (in the sense of McKean) diffusion process obtained by replacing the density $f_{t}$ with $f_{\infty}$
$\mathrm{d}Y_{t}=-\nabla V(Y_{t})\,\mathrm{d}t-(\nabla W*f_{\infty})(Y_{t})\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},\qquad Y_{0}\sim\nu_{0},$ (1.5)
for some initial condition $\nu_{0}\in\mathcal{P}(\mathbb{D})$, are reversible, when started at stationarity, with respect to the same Gibbsian invariant measure $f_{\infty}$, the unique solution of (1.4); see [50, Chapter 4]. The exponentially fast convergence of $f_{t}$ to $f_{\infty}$, suggests that, for arbitrary initial conditions $X_{0}\sim\mu_{0}$ and $Y_{0}\sim\nu_{0}$, we expect that the nonlinear process $X_{t}$ and the linear process $Y_{t}$ are exponentially close in time. This suggests that, in order to study the long-term behavior of solutions to the Mckean SDE and to the corresponding McKean–Vlasov PDE, it is sufficient to consider the linearized process $Y_{t}$. This fact was already used implicitly in [51] when applying the martingale estimating function methodology [36] to estimate parameters in the drift as well as the diffusion coefficient in equation (1.1). In particular, it was shown in [51] that, in order to infer parameters in the mean field dynamics, it is sufficient to estimate the eigenpairs of the generator of the linearized dynamics (1.5). This linearization methodology was also used in [22] when studying the joint diffusive-mean field limit of the interacting particle system (1.1) with periodic coefficients. In particular, it was shown that in order to study the diffusive limit of the mean field dynamics, it is sufficient to prove a central limit theorem and, based on this, an invariance principle for the linearized dynamics. Consequently, the formula for the covariance matrix of the limiting Brownian motion requires the solution of the Poisson equation for the linearized process, and the computation of an average with respect to the Gibbs measure of the process on the torus. We reiterate that the nonlinear and linear processes are reversible with respect to the same Gibbs measure.
In this paper we build on the ideas presented in [22, 51] in order to develop the linearization of the McKean SDE as a systematic tool for studying the mean field limit of weakly interacting diffusions. More specifically, we show that, when confronted with tasks that require the analysis of the long-time behaviour of the dynanics, the nonlinear process $X_{t}$ can be “replaced” by the linear process $Y_{t}$, whenever the confining and interaction potentials are such that uniform propagation of chaos holds. We obtain quantitative estimates on the distance between the two processes both in relative entropy and Wasserstein distance, considering both the whole space and the torus as state spaces. Moreover, in the former setting, we also show strong convergence in $L^{2}$ for the paths using a coupling argument. We remark that the problem of linearizing nonlinear Fokker–Planck equations on the whole space has also been studied in [53], where the authors propose a different method to obtain a linearized process which uses the natural tangent bundle over the space of probabilities with the corresponding gradient operator. Our analysis, instead, is mostly based on entropy estimates between solutions of Fokker–Planck equations presented in [26, 39], based on the important paper [8]. We then combine these estimates with a logarithmic Sobolev inequality for the Gibbs measure and for the nonlinear and linear semigroups. In particular, we have that, for $X_{t}\sim\mu_{t}$ and $Y_{t}\sim\nu_{t}$,
$\mathcal{H}(\mu_{t}|\nu_{t})+\frac{1}{2\beta}\int_{0}^{t}\mathcal{I}(\mu_{s}|\nu_{s})\,\mathrm{d}s\leq\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta}{2}\int_{0}^{t}\int_{\mathbb{D}}\left\|(\nabla W*(\mu_{s}-\mu_{\infty}))(x)\right\|^{2}\mu_{s}(x)\,\mathrm{d}x\,\mathrm{d}s,$ (1.6)
where $\mathcal{H}$, $\mathcal{I}$ denote the relative entropy and relative Fisher information, respectively. This, together with the LSI, implies
$\mathcal{H}(\mu_{t}|\nu_{t})+\frac{1}{2\beta\lambda}\int_{0}^{t}\mathcal{H}(\mu_{s}|\nu_{s})\,\mathrm{d}s\leq\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta}{2}\int_{0}^{t}\int_{\mathbb{D}}\left\|(\nabla W*(\mu_{s}-\mu_{\infty}))(x)\right\|^{2}\mu_{s}(x)\,\mathrm{d}x\,\mathrm{d}s,$ (1.7)
for some appropriate constant $\lambda>0$. The details can be found in Section 2. The fact that the laws $\mu_{t}$ and $\nu_{t}$ are exponentially close follows from this estimate.
We consider the problem both in $\mathbb{R}^{d}$ and in the torus $\mathbb{T}^{d}$. In the latter case, we extend some of the results obtained in [14, 33] on convergence to equilibrium and the LSI to the case where a confining potential in present, in addition to the interaction potential. This is one of the novelties of the present work. Then, we apply our methodology to propose a maximum likelihood estimator (MLE) for the mean field SDE (1.2) which is based on the likelihood of the linearized process $Y_{t}$, and we prove its asymptotic unbiasedness in the limit of infinite observations. Finally, we also revisit the combined diffusive-mean field limit for interacting diffusions in periodic confining and interaction potentials that was studied in [22], and we calculate the variance of the limiting Brownian motion using the linearized process.
The main contributions of this work can be summarized as follows.
  • We prove that, under assumptions on the drift and confining potential that ensure uniform propagation of chaos, the McKean SDE on $\mathbb{R}^{d}$ is exponentially close in time to the linearized process, both in relative entropy and in the $2$-Wasserstein distance.
  • We obtain a similar result for the McKean SDE on the torus in relative entropy. As a by-product, we extend some of the known results on the McKean SDE on the torus to the case where an arbitrary confining potential is present.
  • We use the linearized process (1.5) to propose a “linearized” MLE estimator for the nonlinear McKean process which is asymptotically unbiased. In addition, we study the joint diffusive-mean field limit for weakly interacting diffusions in periodic confining and interaction potentials. Our theoretical results are supplemented by numerical simulations.
The remainder of the paper is organized as follows. In Section 2 we show the convergence analysis for the law of the linearized McKean SDE, which we perform both in the whole space in Section 2.2 and on the torus in Section 2.3. Then, in Section 3 we present two applications of the linearized process: maximum likelihood estimation in Section 3.1 and computation of the asymptotic variance in a central limit theorem in Section 3.2. Finally, in Section 4 we draw our conclusions and outline future developments of the linearization procedure.
2 The linearized diffusion process
2.1 Preliminaries
We consider the McKean SDE (1.2) on the domain $\mathbb{D}$ (either $\mathbb{R}^{d}$ or $\mathbb{T}^{d}$)
$\mathrm{d}X_{t}=-\nabla V(X_{t})\,\mathrm{d}t-(\nabla W*f_{t})(X_{t})\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},$ (2.1)
and its linearization (1.5)
$\mathrm{d}Y_{t}=-\nabla V(Y_{t})\,\mathrm{d}t-(\nabla W*f_{\infty})(Y_{t})\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t}.$ (2.2)
We remark that to define the linearized process $Y_{t}$ we need the following main hypothesis.
Assumption 2.1.
The integral equation (1.4) has a unique solution $f_{\infty}$ and therefore the McKean SDE (1.2) has a unique invariant measure $\mathrm{d}\mu_{\infty}(x)=f_{\infty}(x)\,\mathrm{d}x$.
Let now $\mu_{t}$ and $\nu_{t}$ be the law of the processes $X_{t}$ and $Y_{t}$, respectively, and denote their densities with respect to the Lebesgue measure by $f_{t}$ and $g_{t}$, which satisfy the Fokker–Planck equations
$\frac{\partial f_{t}}{\partial t}$ $=\nabla\cdot(\nabla Vf_{t}+(\nabla W*f_{t})f_{t})+\beta^{-1}\Delta f_{t},$ (2.3)
$\frac{\partial g_{t}}{\partial t}$ $=\nabla\cdot(\nabla Vg_{t}+(\nabla W*f_{\infty})g_{t})+\beta^{-1}\Delta g_{t},$
together with the appropriate initial conditions. Note that, by construction, the invariant measures coincide, that is, $\nu_{\infty}=\mu_{\infty}$ and $g_{\infty}=f_{\infty}$, and the densities are given by the integral equation (1.4). In the following sections, we quantify the distance between processes $X_{t}$ and $Y_{t}$, focusing first on the case where the domain $\mathbb{D}$ is the whole space $\mathbb{R}^{d}$ and then on the multidimensional torus $\mathbb{T}^{d}$. The theoretical results, and in particular the main assumptions that guarantee, e.g., the uniqueness of the invariant measure, are specific for each case and therefore are discussed in each section separately.
Remark 2.2.
The situation is more complicated when we do not have uniqueness of the invariant measure. However, Assumption 2.1 is not a strong limitation of the methodology introduced in this work. In fact, in [44] it is shown that when multiple invariant measures can exist, it is still possible to prove exponentially fast convergence to equilibrium locally, i.e., as long as the initial distribution is sufficiently close to the stationary state and we are away from the critical temperature, where we only get algebraic convergence. Therefore, we believe that the results presented here can be extended to the case of multiple invariant measures by selecting the “correct” density for the linearization procedure depending on basins of attraction of the different invariant measures. In fact, this idea was already employed in the numerical experiments on inference for the Desai–Zwanzig model in [51].
The relative entropy $\mathcal{H}$ and the relative Fisher information $\mathcal{I}$ between two distributions $\mu$ and $\nu$ in $\mathcal{P}(\mathbb{D})$ such that $\mu$ is absolutely continuous with respect to $\nu$ are
$\mathcal{H}(\mu|\nu)=\int_{\mathbb{D}}\log\left(\frac{\mathrm{d}\mu}{\mathrm{d}\nu}(x)\right)\,\mathrm{d}\mu(x),\qquad\mathcal{I}(\mu|\nu)=\int_{\mathbb{D}}\left|\nabla\log\left(\frac{\mathrm{d}\mu}{\mathrm{d}\nu}(x)\right)\right|^{2}\,\mathrm{d}\mu(x).$ (2.4)
Moreover, if $\mu$ and $\nu$ admit densities $f$ and $g$, respectively, with respect to the Lebesgue measure, then the relative entropy and Fisher information can be rewritten as
$\mathcal{H}(\mu|\nu)=\int_{\mathbb{D}}\log\left(\frac{f(x)}{g(x)}\right)f(x)\,\mathrm{d}x,\qquad\mathcal{I}(\mu|\nu)=\int_{\mathbb{D}}\left|\nabla\log\left(\frac{f(x)}{g(x)}\right)\right|^{2}f(x)\,\mathrm{d}x.$ (2.5)
We finally recall that, following the notation in [43], $\nu$ satisfies an LSI with constant $\lambda>0$ if for all $\mu\in\mathcal{P}(\mathbb{D})$ such that $\mu$ is absolutely continuous with respect to $\nu$ it holds
$\mathcal{H}(\mu|\nu)\leq\frac{\lambda}{4}\mathcal{I}(\mu|\nu).$ (2.6)
2.2 Analysis on $\mathbb{R}^{d}$
We now fix the domain $\mathbb{D}$ to be the whole space $\mathbb{R}^{d}$ and we consider the setting of [41], which guarantees the well-posedness of the McKean SDE and of the corresponding McKean–Vlasov PDE and the uniqueness of the invariant measure due to the convexity of the confining and interaction potentials. In particular, we work under the following assumptions.
Assumption 2.3.
The initial distributions $\mu_{0}$ and $\nu_{0}$ have bounded moments of all orders, and the potentials $V,W\colon\mathbb{R}^{d}\to\mathbb{R}$ satisfy:
  1. (i)
    $V,\,W\in\mathcal{C}^{2}(\mathbb{R}^{d})$ are polynomially bounded along with their derivatives;
  2. (ii)
    $W$ is even;
  3. (iii)
    $V$ and $W$ are uniformly convex with constants $\alpha>0$ and $\gamma\geq 0$, respectively, which means that for all $x,y\in\mathbb{R}^{d}$
    $(\nabla V(x)-\nabla V(y))^{\top}(x-y)$ $\geq\alpha\left\lvert x-y\right\rvert^{2},$ (2.7)
    $(\nabla W(x)-\nabla W(y))^{\top}(x-y)$ $\geq\gamma\left\lvert x-y\right\rvert^{2}.$
Then, from [41, Theorem 1.3] it follows that the density $f_{t}$ converges to $f_{\infty}$ in $L^{1}$ exponentially fast, i.e., there exists a constant $K>0$ such that for all $t\geq 0$
$\left\|f_{t}-f_{\infty}\right\|_{L^{1}(\mathbb{R}^{d})}\leq Ke^{-\frac{1}{2}\alpha t}.$ (2.8)
Moreover, since $\mu_{0}$ has finite moments of any order, then so does the law of the process uniformly in time. Therefore, we have the following bound.
Lemma 2.4.
Under Assumption 2.3, there exists $M>0$ such that for all $t\geq 0$
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left\lvert\nabla W(x-z)\right\rvert^{2}(f_{s}(z)+f_{\infty}(z))f_{s}(x)\,\mathrm{d}z\,\mathrm{d}x\leq M.$ (2.9)
Proof.
Since $\mu_{t}$ and $\mu_{\infty}$ have bounded moments of any order uniformly in time by [41], the desired result follows from the fact that $\nabla W$ is polynomially bounded. ∎
In the following lemma we show that convolving the interaction potential with a probability density function preserves the convexity property.
Lemma 2.5.
Let $W$ satisfy Assumption 2.3(iii), then, for all probability density functions $f$ of measures in $\mathcal{P}(\mathbb{R}^{d})$, the function $W*f$ is uniformly convex with constant $\gamma\geq 0$, i.e., it holds for all $x,y\in\mathbb{R}^{d}$
$((\nabla W*f)(x)-(\nabla W*f)(y))^{\top}(x-y)\geq\gamma\left\lvert x-y\right\rvert^{2}.$ (2.10)
Proof.
By definition of convolution, we have
$((\nabla W*f)(x)-(\nabla W*f)(y))^{\top}(x-y)$ (2.11)
$\qquad=\left(\int_{\mathbb{R}^{d}}(\nabla W(x-z)-\nabla W(y-z))f(z)\,\mathrm{d}z\right)^{\top}(x-y)$
$\qquad=\int_{\mathbb{R}^{d}}(\nabla W(x-z)-\nabla W(y-z))^{\top}((x-z)-(y-z))f(z)\,\mathrm{d}z,$
which due to Assumption 2.3(iii) implies
$((\nabla W*f)(x)-(\nabla W*f)(y))^{\top}(x-y)\geq\int_{\mathbb{R}^{d}}\gamma\left\lvert x-y\right\rvert^{2}f(z)\,\mathrm{d}z=\gamma\left\lvert x-y\right\rvert^{2},$ (2.12)
which is the desired result. ∎
2.2.1 Convergence of the linearized process
In the classical work [41] as well in the interesting recent paper [43], it is shown that, provided that the initial condition satisfies an LSI, and under our convexity assumptions on the confining and interaction potentials given in Assumption 2.3, then so does the solution of the McKean–Vlasov PDE (1.3), with an LSI constant that can be explicitly computed. In the following lemma we show that the same conclusion holds for the linearized McKean–Vlasov PDE, with the same LSI constant.
Lemma 2.6.
Let Assumption 2.3 be satisfied and assume that $\mu_{0}$ and $\nu_{0}$ satisfy an LSI with constant $\lambda_{0}>0$. Then $\mu_{t}$ and $\nu_{t}$ satisfy an LSI with constant
$\lambda_{t}=\lambda_{0}e^{-2(\alpha+\gamma)t}+\frac{1}{2\beta(\alpha+\gamma)}\left(1-e^{-2(\alpha+\gamma)t}\right).$ (2.13)
Moreover, $\mu_{t}$ and $\nu_{t}$ satisfy a time-uniform LSI with constant
$\Lambda=\max\left\{\lambda_{0},\frac{1}{2\beta(\alpha+\gamma)}\right\}.$ (2.14)
Proof.
The desired result follows directly from [43, Proposition 1.1] and Lemma 2.5, and the uniform in time constant is obtained by noticing that $\lambda_{t}$ is a convex combination of $\lambda_{0}$ and $1/(2\beta(\alpha+\gamma))$. ∎
From the LSI we can now deduce the exponentially fast convergence in relative entropy of the law of the linearized process $\nu_{t}$ to the law of the nonlinear process $\mu_{t}$.
Theorem 2.7.
Let the assumptions of Lemma 2.6 be satisfied. Then, it holds
$\mathcal{H}(\mu_{t}|\nu_{t})\leq\rho_{\Lambda}(t)\coloneqq\begin{cases}\left(\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta^{2}MK\Lambda}{4-\alpha\beta\Lambda}\right)e^{-\frac{1}{2}\alpha t},&\text{if }\Lambda<\frac{4}{\alpha\beta},\\ \left(\mathcal{H}(\mu_{0}|\nu_{0})+\frac{1}{2}\beta MKt\right)e^{-\frac{2}{\beta\Lambda}t},&\text{if }\Lambda=\frac{4}{\alpha\beta},\\ \left(\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta^{2}MK\Lambda}{\alpha\beta\Lambda-4}\right)e^{-\frac{2}{\beta\Lambda}t},&\text{if }\Lambda>\frac{4}{\alpha\beta}.\end{cases}$ (2.15)
Proof.
By [39, Lemma 3.1] we have the estimate
$\mathcal{H}(\mu_{t}|\nu_{t})+\frac{1}{2\beta}\int_{0}^{t}\mathcal{I}(\mu_{s}|\nu_{s})\,\mathrm{d}s\leq\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta}{2}\int_{0}^{t}\int_{\mathbb{R}^{d}}\left\lvert(\nabla W*(f_{s}-f_{\infty}))(x)\right\rvert^{2}f_{s}(x)\,\mathrm{d}x\,\mathrm{d}s,$ (2.16)
which due to Lemma 2.6 implies
$\mathcal{H}(\mu_{t}|\nu_{t})+\frac{2}{\beta\Lambda}\int_{0}^{t}\mathcal{H}(\mu_{s}|\nu_{s})\,\mathrm{d}s\leq\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta}{2}\int_{0}^{t}\int_{\mathbb{R}^{d}}\left\lvert(\nabla W*(f_{s}-f_{\infty}))(x)\right\rvert^{2}f_{s}(x)\,\mathrm{d}x\,\mathrm{d}s.$ (2.17)
By Cauchy–Schwarz inequality we have
$\left\lvert(\nabla W*(f_{s}-f_{\infty}))(x)\right\rvert^{2}\leq\left(\int_{\mathbb{R}^{d}}\left\lvert\nabla W(x-z)\right\rvert^{2}(f_{s}(z)+f_{\infty}(z))\,\mathrm{d}z\right)\left\|f_{s}-f_{\infty}\right\|_{L^{1}(\mathbb{R}^{d})},$ (2.18)
and due to the bound (2.9) and the convergence to the invariant measure given in (2.8) we obtain
$\int_{\mathbb{R}^{d}}\left\lvert(\nabla W*(f_{s}-f_{\infty}))(x)\right\rvert^{2}f_{s}(x)\,\mathrm{d}x\leq M\left\|f_{s}-f_{\infty}\right\|_{L^{1}(\mathbb{R}^{d})}\leq MKe^{-\frac{1}{2}\alpha s}.$ (2.19)
Therefore, estimate (2.17) becomes
$\mathcal{H}(\mu_{t}|\nu_{t})\leq\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta MK}{2}\int_{0}^{t}e^{-\frac{1}{2}\alpha s}\,\mathrm{d}s-\frac{2}{\beta\Lambda}\int_{0}^{t}\mathcal{H}(\mu_{s}|\nu_{s})\,\mathrm{d}s,$ (2.20)
which, by [39, Lemma A.1], implies
$\mathcal{H}(\mu_{t}|\nu_{t})$ $\leq\mathcal{H}(\mu_{0}|\nu_{0})e^{-\frac{2}{\beta\Lambda}t}+\frac{\beta MK}{2}\int_{0}^{t}e^{-\frac{2}{\beta\Lambda}(t-s)}e^{-\frac{1}{2}\alpha s}\,\mathrm{d}s,$ (2.21)
$=\mathcal{H}(\mu_{0}|\nu_{0})e^{-\frac{2}{\beta\Lambda}t}+\frac{\beta MK}{2}e^{-\frac{2}{\beta\Lambda}t}\int_{0}^{t}e^{\frac{4-\alpha\beta\Lambda}{2\beta\Lambda}s}\,\mathrm{d}s$
Then, if $\Lambda=4/(\alpha\beta)$ we deduce
$\mathcal{H}(\mu_{t}|\nu_{t})\leq\left(\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta MK}{2}t\right)e^{-\frac{2}{\beta\Lambda}t}.$ (2.22)
On the other hand, if $\Lambda\neq 4/(\alpha\beta)$ we have
$\mathcal{H}(\mu_{t}|\nu_{t})\leq\mathcal{H}(\mu_{0}|\nu_{0})e^{-\frac{2}{\beta\Lambda}t}+\frac{\beta^{2}MK\Lambda}{4-\alpha\beta\Lambda}\left(e^{-\frac{1}{2}\alpha t}-e^{-\frac{2}{\beta\Lambda}t}\right),$ (2.23)
and the desired result follows. ∎
The following convergence in Wasserstein distance is a direct consequence of the Talagrand’s transportation inequality.
Corollary 2.8.
Let the assumptions of Theorem 2.7 be satisfied. Then, it holds
$\mathcal{W}_{2}(\mu_{t},\nu_{t})\leq\sqrt{\Lambda\rho_{\Lambda}(t)},$ (2.24)
where $\mathcal{W}_{2}$ denotes the 2-Wasserstein distance and $\rho_{\Lambda}(t)$ is defined in Theorem 2.7.
Proof.
The desired result follows from Theorem 2.7 using [48, Theorem 1]. ∎
Finally, we also have strong convergence of the paths $X_{t}$ and $Y_{t}$ in $L^{2}$.
Theorem 2.9.
Let Assumption 2.3 be satisfied. Then, it holds
$\operatorname{\mathbb{E}}[\left\|X_{t}-Y_{t}\right\|^{2}]\leq\left(\operatorname{\mathbb{E}}[\left\|X_{0}-Y_{0}\right\|^{2}]+\frac{8MK}{(3\alpha+4\gamma)^{2}}\right)e^{-\frac{1}{2}\alpha t}.$ (2.25)
Proof.
Combining equations (1.2) and (1.5), we have
$\frac{\mathrm{d}}{\mathrm{d}t}(X_{t}-Y_{t})=-(\nabla V(X_{t})-\nabla V(Y_{t}))-((\nabla W*f_{t})(X_{t})-(\nabla W*f_{\infty})(Y_{t})),$ (2.26)
which, multiplying by $(X_{t}-Y_{t})$, implies
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\left\lvert X_{t}-Y_{t}\right\rvert^{2}$ $=-(\nabla V(X_{t})-\nabla V(Y_{t}))^{\top}(X_{t}-Y_{t})$ (2.27)
$\quad-((\nabla W*f_{t})(X_{t})-(\nabla W*f_{\infty})(Y_{t}))^{\top}(X_{t}-Y_{t})$
$\eqqcolon I_{1}+I_{2}.$
First, by the uniform convexity of $V$ we get
$I_{1}\leq-\alpha\left\lvert X_{t}-Y_{t}\right\rvert^{2},$ (2.28)
and then we rewrite $I_{2}$ as
$I_{2}$ $=-((\nabla W*f_{t})(X_{t})-(\nabla W*f_{\infty})(X_{t}))^{\top}(X_{t}-Y_{t})$ (2.29)
$\quad-((\nabla W*f_{\infty})(X_{t})-(\nabla W*f_{\infty})(Y_{t}))^{\top}(X_{t}-Y_{t})$
$\eqqcolon I_{2,1}+I_{2,2},$
where Lemma 2.5 gives
$I_{2,2}\leq-\gamma\left\lvert X_{t}-Y_{t}\right\rvert^{2}.$ (2.30)
Using Young’s inequality with $\delta>0$ we have
$I_{2,1}\leq\frac{1}{2\delta}\left\lvert(\nabla W*f_{t})(X_{t})-(\nabla W*f_{\infty})(X_{t})\right\rvert^{2}+\frac{\delta}{2}\left\lvert X_{t}-Y_{t}\right\rvert^{2},$ (2.31)
and by Cauchy–Schwarz inequality we obtain
$I_{2,1}\leq\frac{\delta}{2}\left\lvert X_{t}-Y_{t}\right\rvert^{2}+\frac{1}{2\delta}\left(\int_{\mathbb{R}^{d}}\left\lvert\nabla W(X_{t}-z)\right\rvert^{2}(f_{t}(z)+f_{\infty}(z))\,\mathrm{d}z\right)\left\|f_{t}-f_{\infty}\right\|_{L^{1}(\mathbb{R}^{d})}.$ (2.32)
Collecting bounds (2.28), (2.30), (2.32) and by equations (2.8) and (2.9) we have
$\frac{\mathrm{d}}{\mathrm{d}t}\operatorname{\mathbb{E}}[\left\lvert X_{t}-Y_{t}\right\rvert^{2}]\leq-2\left(\alpha+\gamma-\frac{\delta}{2}\right)\operatorname{\mathbb{E}}[\left\lvert X_{t}-Y_{t}\right\rvert^{2}]+\frac{MK}{2\delta}e^{-\frac{1}{2}\alpha t},$ (2.33)
which due to Grönwall’s inequality implies
$\operatorname{\mathbb{E}}[\left\lvert X_{t}-Y_{t}\right\rvert^{2}]$ $\leq\operatorname{\mathbb{E}}[\left\lvert X_{0}-Y_{0}\right\rvert^{2}]e^{-\left(2\alpha+2\gamma-\delta\right)t}+\frac{MK}{2\delta}\int_{0}^{t}e^{-\left(2\alpha+2\gamma-\delta\right)(t-s)}e^{-\frac{1}{2}\alpha s}\,\mathrm{d}s$ (2.34)
$=\operatorname{\mathbb{E}}[\left\lvert X_{0}-Y_{0}\right\rvert^{2}]e^{-\left(2\alpha+2\gamma-\delta\right)t}+\frac{MK}{\delta(3\alpha+4\gamma-2\delta)}\left(e^{-\frac{1}{2}\alpha t}-e^{-\left(2\alpha+2\gamma-\delta\right)t}\right).$
Finally, choosing the optimal value $\delta=(3\alpha+4\gamma)/4$ yields the desired result. ∎
Remark 2.10.
If the LSI constant $\lambda_{0}$ of the initial distributions $\mu_{0}$ and $\nu_{0}$ is sufficiently small, namely $\lambda_{0}<4/(\alpha\beta)$, then the rate of convergence in Theorems 2.7 and 2.8 is determined only by the convexity parameter $\alpha$ of the confining potential. This is in agreement with the rate of convergence in $L^{2}$ predicted by Theorem 2.9.
2.3 Analysis on $\mathbb{T}^{d}$
In this section, the domain $\mathbb{D}$ is the multidimensional torus $\mathbb{T}^{d}$ of size length $1$, and we work in the framework below.
Assumption 2.11.
The potentials $V,W\colon\mathbb{T}^{d}\to\mathbb{R}$ satisfy:
  1. (i)
    $V,W\in\mathcal{C}^{2}(\mathbb{T}^{d})$;
  2. (ii)
    $W$ is even;
  3. (iii)
    W is either $H$-stable, meaning that it has nonnegative Fourier modes [55], or is such that $\left\|W\right\|_{L^{\infty}(\mathbb{T}^{d})}<\beta^{-1}$.
Notice that Assumption 2.11(iii) guarantees that Assumption 2.1 is satisfied. In fact, all invariant probability density functions $f_{\infty}$ are also minimizers of the free energy
$\mathcal{E}(f)=\int_{\mathbb{T}^{d}}V(x)f(x)\,\mathrm{d}x+\frac{1}{2}\int_{\mathbb{T}^{d}}\int_{\mathbb{T}^{d}}W(x-y)f(y)f(x)\,\mathrm{d}y\,\mathrm{d}x+\beta^{-1}\int_{\mathbb{T}^{d}}f(x)\log(f(x))\,\mathrm{d}x,$ (2.35)
which is strictly convex under Assumption 2.11(iii). This is shown in [14, Proposition 2.8] without the confining potential, which, however, does not affect the convexity of $\mathcal{E}$ since its corresponding term is linear; see also [3, Section V.B.]. Moreover, the invariant measure $\mu_{\infty}$ satisfies an LSI, as shown in the following result.
Lemma 2.12.
Under Assumption 2.11, the invariant measure $\mathrm{d}\mu_{\infty}(x)=f_{\infty}(x)\,\mathrm{d}x$ satisfies an LSI with constant $\Gamma/(2\pi^{2})$, where
$\Gamma=e^{2\beta\left(\left\|V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|W\right\|_{L^{\infty}(\mathbb{T}^{d})}\right)}.$ (2.36)
Proof.
We follow [33, Lemma 2.4]. From the integral equation (1.4) we deduce that for all $x\in\mathbb{T}^{d}$
$\frac{1}{\Gamma}=e^{-2\beta\left(\left\|V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|W\right\|_{L^{\infty}(\mathbb{T}^{d})}\right)}\leq f_{\infty}(x)\leq e^{2\beta\left(\left\|V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|W\right\|_{L^{\infty}(\mathbb{T}^{d})}\right)}=\Gamma.$ (2.37)
Then, the desired result follows from the Holley–-Stroock perturbation lemma (see, e.g., [33, Lemma 2.3]) and the fact that the uniform distribution on the torus satisfies an LSI (see, e.g., [33, Lemma 2.4]). We remark that the discrepancy with [33, Lemma 2.4] is due to the factor $4$ in the LSI (2.6). ∎
Remark 2.13.
The results presented in the following sections still hold if, instead of having a unique invariant measure, we have multiple stationary states but a unique stable one. This is the case, e.g, of the model in [22, Lemma 1.26] above the phase transition. In fact, for all initial conditions except the unstable invariant measure, the law of the McKean SDE converges to the stable stationary state. It is not clear whether it is possible to have a similar situation (precisely two invariant measures, one stable and one unstable) when the state space is the whole space $\mathbb{R}^{d}$, at least for polynomial interaction and confining potentials [61].
2.3.1 Convergence to equilibrium for the McKean–Vlasov PDE
Before presenting the convergence results for the linearized process, we need the exponentially fast convergence of the measure $\mu_{t}$ to the invariant measure $\mu_{\infty}$. This problem has already been studied when there is no confining potential in, e.g., [17, 14]. In the absence of confining potential, the analysis is simplified because the unique, under Assumption 2.11(iii), invariant measure is the Lebesgue measure. However, here we extend the results to the case of a nonzero confining potential. Given two probabilities $\mu,\nu\in\mathcal{P}(\mathbb{T}^{d})$ with densities $f,g$, we denote their total variation distance by $\mathrm{TV}(\mu,\nu)$. Moreover, we recall that it holds the following equivalence with the $L^{1}(\mathbb{T}^{d})$ distance between the densities
$\mathrm{TV}(\mu,\nu)=\frac{1}{2}\left\|f-g\right\|_{L^{1}(\mathbb{T}^{d})}.$ (2.38)
We first consider the $L^{2}(\mathbb{T}^{d})$ distance between the densities $f_{t}$ and $f_{\infty}$, whose analysis is inspired by the proof of [17, Theorem 2.3], and then move to the convergence in relative entropy, for which we generalize the proof of [14, Proposition 3.1].
Proposition 2.14.
Let Assumption 2.11 be satisfied. Then, it holds
$\left\|f_{t}-f_{\infty}\right\|_{L^{2}(\mathbb{T}^{d})}\leq\left\|f_{0}-f_{\infty}\right\|_{L^{2}(\mathbb{T}^{d})}e^{-\zeta t},$ (2.39)
where
$\zeta=\frac{\pi^{2}}{4\beta}-\beta\left(\left\|\nabla V\right\|_{L^{\infty}(\mathbb{T}^{d})}+(1+\Gamma)\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}\right)^{2},$ (2.40)
and $\Gamma$ is defined in equation (2.36).
Proof.
We rewrite the McKean–Vlasov PDE (1.3) for the difference $f_{t}-f_{\infty}$, using the fact that $f_{\infty}$ solves the stationary Fokker–Planck equation. We then multiply by $f_{t}-f_{\infty}$ and integrate by parts to obtain
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\left\|f_{t}-f_{\infty}\right\|_{L^{2}(\mathbb{T}^{d})}^{2}$ $=\int_{\mathbb{T}^{d}}\nabla\cdot\left((\nabla V+\nabla W*f_{t})f_{t}-(\nabla V+\nabla W*f_{\infty})f_{\infty}\right)(f_{t}-f_{\infty})$ (2.41)
$\quad+\beta^{-1}\int_{\mathbb{T}^{d}}\Delta(f_{t}-f_{\infty})(f_{t}-f_{\infty})$
$=-\int_{\mathbb{T}^{d}}\left((\nabla V+\nabla W*f_{t})f_{t}-(\nabla V+\nabla W*f_{\infty})f_{\infty}\right)\cdot\nabla(f_{t}-f_{\infty})$
$\quad-\beta^{-1}\left\|\nabla(f_{t}-f_{\infty})\right\|_{L^{2}(\mathbb{T}^{d})}^{2}$
$\eqqcolon I_{t}-\beta^{-1}\left\|\nabla(f_{t}-f_{\infty})\right\|_{L^{2}(\mathbb{T}^{d})}^{2}.$
The quantity $I_{t}$ can then be decomposed as
$I_{t}=-\int_{\mathbb{T}^{d}}(\nabla V+\nabla W*f_{t})(f_{t}-f_{\infty})\cdot\nabla(f_{t}-f_{\infty})-\int_{\mathbb{T}^{d}}\nabla W*(f_{t}-f_{\infty})\cdot\nabla(f_{t}-f_{\infty})f_{\infty},$ (2.42)
which, due to Hölder’s and Young’s convolution inequalities, implies
$\left\lvert I_{t}\right\rvert$ $\leq\left(\left\|\nabla V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}\right)\left\|f_{t}-f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}\left\|\nabla(f_{t}-f_{\infty})\right\|_{L^{1}(\mathbb{T}^{d})}$ (2.43)
$\quad+\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}\left\|f_{\infty}\right\|_{L^{\infty}(\mathbb{T}^{d})}\left\|f_{t}-f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}\left\|\nabla(f_{t}-f_{\infty})\right\|_{L^{1}(\mathbb{T}^{d})}.$
Since the $L^{1}(\mathbb{T}^{d})$ norm can be bounded by the $L^{2}(\mathbb{T}^{d})$ norm, we have
$\left\lvert I_{t}\right\rvert\leq\left(\left\|\nabla V\right\|_{L^{\infty}(\mathbb{T}^{d})}+(1+\Gamma)\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}\right)\left\|f_{t}-f_{\infty}\right\|_{L^{2}(\mathbb{T}^{d})}\left\|\nabla(f_{t}-f_{\infty})\right\|_{L^{2}(\mathbb{T}^{d})},$ (2.44)
where $\Gamma$ is defined in equation (2.36), and which, by Young’s inequality with $\delta>0$, gives
$\left\lvert I_{t}\right\rvert$ $\leq\frac{\delta}{2}\left(\left\|\nabla V\right\|_{L^{\infty}(\mathbb{T}^{d})}+(1+\Gamma)\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}\right)^{2}\left\|f_{t}-f_{\infty}\right\|_{L^{2}(\mathbb{T}^{d})}^{2}$ (2.45)
$\quad+\frac{1}{2\delta}\left\|\nabla(f_{t}-f_{\infty})\right\|_{L^{2}(\mathbb{T}^{d})}^{2}.$
Using the bound (2.45) in equation (2.41) and setting $\delta=\beta$, we obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\left\|f_{t}-f_{\infty}\right\|_{L^{2}(\mathbb{T}^{d})}^{2}$ $\leq-\beta^{-1}\left\|\nabla(f_{t}-f_{\infty})\right\|_{L^{2}(\mathbb{T}^{d})}^{2}$ (2.46)
$\quad+\beta\left(\left\|\nabla V\right\|_{L^{\infty}(\mathbb{T}^{d})}+(1+\Gamma)\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}\right)^{2}\left\|f_{t}-f_{\infty}\right\|_{L^{2}(\mathbb{T}^{d})}^{2},$
and, since $f_{t}-f_{\infty}$ has zero mean, applying Poincaré inequality with the optimal constant $4/\pi^{2}$ implies
$\frac{\mathrm{d}}{\mathrm{d}t}\left\|f_{t}-f_{\infty}\right\|_{L^{2}(\mathbb{T}^{d})}^{2}\leq-\zeta\left\|f_{t}-f_{\infty}\right\|_{L^{2}(\mathbb{T}^{d})}^{2},$ (2.47)
where $\zeta$ is defined in equation (2.40). A proof of the Poincaré inequality on the torus in one dimension can be found, e.g., in [2, Proposition 4.5.5], while the result in the multidimensional setting follows from the tensorization property. Finally, applying Grönwall’s inequality completes the proof. ∎
Proposition 2.15.
Let Assumption 2.11 be satisfied. Then, it holds
$\mathcal{H}(f_{t}|f_{\infty})\leq\mathcal{H}(f_{0}|f_{\infty})e^{-\eta t},$ (2.48)
where
$\eta=8\left(\frac{\pi^{2}}{\beta\Gamma}-\left\|\Delta W\right\|_{L^{\infty}(\mathbb{T}^{d})}-\beta\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}(\left\|\nabla V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})})\right),$ (2.49)
and $\Gamma$ is defined in equation (2.36).
Proof.
By differentiating the relative entropy and since $f_{t}$ solves the McKean–Vlasov PDE (1.3), we have
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{H}(f_{t}|f_{\infty})$ $=\int_{\mathbb{T}^{d}}\frac{\partial}{\partial t}f_{t}\left(\log f_{t}-\log f_{\infty}\right)+\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{T}^{d}}f_{t}$ (2.50)
$=\int_{\mathbb{T}^{d}}\left(\nabla\cdot(\nabla Vf_{t}+(\nabla W*f_{t})f_{t})+\beta^{-1}\Delta f_{t}\right)\left(\log f_{t}-\log f_{\infty}\right),$
which, integrating by parts, gives
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{H}(f_{t}|f_{\infty})$ $=\int_{\mathbb{T}^{d}}(\Delta V+\Delta W*f_{t})f_{t}+\int_{\mathbb{T}^{d}}\nabla\log f_{\infty}\cdot(\nabla V+\nabla W*f_{t})f_{t}$ (2.51)
$\quad-\beta^{-1}\int_{\mathbb{T}^{d}}\left\lvert\nabla\log f_{t}\right\rvert^{2}f_{t}+\beta^{-1}\int_{\mathbb{T}^{d}}\nabla\log f_{\infty}\cdot\nabla\log f_{t}f_{t}.$
By adding and subtracting the Fisher information $\beta^{-1}\mathcal{I}(f_{t}|f_{\infty})$ we then get
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{H}(f_{t}|f_{\infty})$ $=\int_{\mathbb{T}^{d}}(\Delta V+\Delta W*f_{t})f_{t}+\int_{\mathbb{T}^{d}}\nabla\log f_{\infty}\cdot(\nabla V+\nabla W*f_{t})f_{t}$ (2.52)
$\quad-\beta^{-1}\mathcal{I}(f_{t}|f_{\infty})+\beta^{-1}\int_{\mathbb{T}^{d}}\left\lvert\nabla\log f_{\infty}\right\rvert^{2}f_{t}-\beta^{-1}\int_{\mathbb{T}^{d}}\nabla\log f_{\infty}\cdot\nabla\log f_{t}f_{t},$
which, integrating by parts and since $f_{\infty}$ solves the integral equation (1.4), implies
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{H}(f_{t}|f_{\infty})$ $=-\beta^{-1}\mathcal{I}(f_{t}|f_{\infty})+\int_{\mathbb{T}^{d}}(\Delta V+\Delta W*f_{t})f_{t}-\int_{\mathbb{T}^{d}}(\Delta V+\Delta W*f_{\infty})f_{t}$ (2.53)
$\quad+\int_{\mathbb{T}^{d}}\nabla\log f_{\infty}\cdot(\nabla V+\nabla W*f_{t})f_{t}-\int_{\mathbb{T}^{d}}\nabla\log f_{\infty}\cdot(\nabla V+\nabla W*f_{\infty})f_{t}$
$=-\beta^{-1}\mathcal{I}(f_{t}|f_{\infty})+\int_{\mathbb{T}^{d}}\Delta W*(f_{t}-f_{\infty})f_{t}+\int_{\mathbb{T}^{d}}\nabla\log f_{\infty}\cdot\nabla W*(f_{t}-f_{\infty})f_{t}.$
Let us notice that, due to an integration by parts, it holds
$\int_{\mathbb{T}^{d}}\Delta W*(f_{t}-f_{\infty})f_{\infty}+\int_{\mathbb{T}^{d}}\nabla\log f_{\infty}\cdot\nabla W*(f_{t}-f_{\infty})f_{\infty}=0,$ (2.54)
and therefore adding this quantity to the right-hand side of equation (2.53) we obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{H}(f_{t}|f_{\infty})$ $=-\beta^{-1}\mathcal{I}(f_{t}|f_{\infty})+\int_{\mathbb{T}^{d}}\Delta W*(f_{t}-f_{\infty})(f_{t}-f_{\infty})$ (2.55)
$\quad+\int_{\mathbb{T}^{d}}\nabla\log f_{\infty}\cdot\nabla W*(f_{t}-f_{\infty})(f_{t}-f_{\infty})$
$\eqqcolon-\beta^{-1}\mathcal{I}(f_{t}|f_{\infty})+I^{(1)}_{t}+I^{(2)}_{t}.$
We now need to bound the terms in the right-hand side. First, by Hölder’s and Young’s convolution inequalities, we have
$\left\lvert I^{(1)}_{t}\right\rvert$ $\leq\left\|\Delta W\right\|_{L^{\infty}(\mathbb{T}^{d})}\left\|f_{t}-f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}^{2},$ (2.56)
$\left\lvert I^{(2)}_{t}\right\rvert$ $\leq\beta\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}(\left\|\nabla V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})})\left\|f_{t}-f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}^{2},$
and then, due to equation (2.38) and the Csiszár–Kullback–Pinsker (CKP) inequality (see, e.g., [10]), we get
$\left\|f_{t}-f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}^{2}\leq 8\mathcal{H}(f_{t}|f_{\infty}).$ (2.57)
Finally, Lemma 2.12 yields
$\mathcal{I}(f_{t}|f_{\infty})\geq\frac{8\pi^{2}}{\Gamma}\mathcal{H}(f_{t}|f_{\infty}),$ (2.58)
which implies
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{H}(f_{t}|f_{\infty})\leq-\eta\mathcal{H}(f_{t}|f_{\infty}),$ (2.59)
where $\eta$ is defined in equation (2.49). The desired result is then obtained applying Grönwall’s inequality. ∎
Remark 2.16.
In Propositions 2.14 and 2.15, the exponents $\zeta$ and $\eta$ are always strictly positive in the high temperature regime, i.e., when $\beta$ is sufficiently large. Moreover, from the proof of [13, Proposition 3.1] it follows that we can replace $\left\|\Delta W\right\|_{L^{\infty}(\mathbb{T}^{d})}$ in (2.49) with $\left\|\Delta W_{u}\right\|_{L^{\infty}(\mathbb{T}^{d})}$, where $W_{u}$ denotes the non-$H$-stable part of $W$ and is defined in [13, Definition 2.1]. However, the same argument cannot be applied to $\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}$.
The corollary below is a straightforward consequence of Propositions 2.14 and 2.15, and quantifies the rate of convergence in $L^{1}(\mathbb{T}^{d})$.
Corollary 2.17.
Let Assumption 2.11 be satisfied. Then, it holds for $i=1,2$
$\left\|f_{t}-f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}\leq c_{i}e^{-a_{i}t},$ (2.60)
with $a_{1}=\zeta$, $a_{2}=\eta/2$, where $\zeta,\eta$ are defined in (2.40) and (2.49), respectively, and
$c_{1}=\left\|f_{0}-f_{\infty}\right\|_{L^{2}(\mathbb{T}^{d})},\qquad c_{2}=2\sqrt{2\mathcal{H}(f_{0}|f_{\infty})}.$ (2.61)
Proof.
The desired result follows applying Propositions 2.14 and 2.15, the CKP inequality (2.57), and bounding the $L^{1}(\mathbb{T}^{d})$ norm with the $L^{2}(\mathbb{T}^{d})$ one. ∎
2.3.2 Convergence of the linearized process
Similarly to the analysis for the whole space, we first need to prove that the laws of the processes satisfy an LSI uniformly in time. In the lemma below, whose argument is inspired from the proof of [39, Lemma 5.1], we show that this holds true as long as the density of the initial condition is bounded away from zero.
Lemma 2.18.
Let Assumption 2.11 be satisfied, and assume that there exists a constant $\kappa>1$ such that for all $x\in\mathbb{T}^{d}$
$\frac{1}{\kappa}\leq f_{0}(x)\leq\kappa\qquad\text{and}\qquad\frac{1}{\kappa}\leq g_{0}(x)\leq\kappa,$ (2.62)
where $f_{0}$ and $g_{0}$ are the densities with respect to the Lebesgue measure of the initial distributions $\mu_{0}$ and $\nu_{0}$. Then, $\nu_{t}$ satisfies a time-uniform LSI with constant
$\Xi=\frac{\kappa\Gamma^{2}}{2\pi^{2}},$ (2.63)
where $\Gamma$ is defined in (2.36). Moreover, if $a_{i}>0$ and $C_{i}/a_{i}<\mathbb{W}(1)$ for either $i=1$ or $i=2$, with $a_{i}$ and $C_{i}$ given in Corollary 2.17 and equation (2.77) below, and where $\mathbb{W}$ denotes the Lambert function, then $\mu_{t}$ satisfies a time-uniform LSI with constant
$\widetilde{\Xi}_{i}=\frac{\kappa\Gamma^{2}}{2\pi^{2}(1-(C_{i}/a_{i})e^{C_{i}/a_{i}})}.$ (2.64)
Proof.
Let $\varphi_{t}$ and $\psi_{t}$ be the Radon–Nikodym derivative of $\nu_{t}$ and $\mu_{t}$ with respect to $\mu_{\infty}=\nu_{\infty}$, i.e.,
$\varphi_{t}=\frac{\mathrm{d}\nu_{t}}{\mathrm{d}\mu_{\infty}}=\frac{g_{t}}{f_{\infty}}\qquad\text{and}\qquad\psi_{t}=\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\mu_{\infty}}=\frac{f_{t}}{f_{\infty}},$ (2.65)
and note that, since $f_{t},g_{t}$ and $f_{\infty}$ solve the time-dependent and stationary Fokker–Planck equations (2.3), respectively, $\varphi_{t}$ solves
$\frac{\partial\varphi_{t}}{\partial t}=\beta^{-1}\Delta\varphi_{t}-(\nabla V+\nabla W*f_{\infty})\cdot\nabla\varphi_{t},$ (2.66)
and $\psi_{t}$ satisfies
$\frac{\partial\psi_{t}}{\partial t}$ $=\beta^{-1}\Delta\psi_{t}+\Delta W*((\psi_{t}-1)f_{\infty})\psi_{t}-\beta(\nabla V+\nabla W*f_{\infty})\cdot\nabla W*((\psi_{t}-1)f_{\infty})\psi_{t}$ (2.67)
$\quad-(\nabla V+\nabla W*f_{\infty})\cdot\nabla\psi_{t}+\nabla W*((\psi_{t}-1)f_{\infty})\cdot\nabla\psi_{t}.$
We now follow the proof of [39, Lemma 5.1] that is based on the argument presented in [33, Theorem 2]; see also the proof of [45, Theorem 5]. Fix $T>0$ and $x\in\mathbb{T}^{d}$, and let $(Z_{t}^{\nu})_{t\in[0,T]}$ be the solution of the SDE
$\mathrm{d}Z_{t}^{\nu}=-(\nabla V(Z_{t}^{\nu})+\nabla W*f_{\infty}(Z_{t}^{\nu}))\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},$ (2.68)
with initial condition $Z_{0}^{\nu}=x$. Using equation (2.66) and applying Itô lemma to the function $\varphi_{T-t}$, we obtain
$\mathrm{d}\varphi_{T-t}(Z_{t}^{\nu})=\sqrt{2\beta^{-1}}\nabla\varphi_{T-t}(Z_{t}^{\nu})\cdot\mathrm{d}B_{t},$ (2.69)
which, integrating, taking the expectation, and noting that the Itô integral has expectation zero, gives
$\operatorname{\mathbb{E}}[\varphi_{T-t}(Z_{t}^{\nu})]=\varphi_{T}(x).$ (2.70)
Then, due to hypothesis (2.62), choosing $t=T$ yields for all $T>0$
$\frac{1}{\kappa\Gamma}\leq\varphi_{T}(x)\leq\kappa\Gamma.$ (2.71)
The argument is similar, but more complex, for the measure $\mu_{t}$. Let $(Z_{t}^{\mu})_{t\in[0,T]}$ be the solution of the SDE
$\mathrm{d}Z_{t}^{\mu}=-(\nabla V(Z_{t}^{\mu})+\nabla W*f_{\infty}(Z_{t}^{\mu}))\,\mathrm{d}t+\nabla W*((\psi_{T-t}-1)f_{\infty})(Z_{t}^{\mu})\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},$ (2.72)
with initial condition $Z_{0}^{\mu}=x$. Using equation (2.67) and applying Itô lemma to the function $\psi_{T-t}$, we obtain
$\mathrm{d}\psi_{T-t}(Z_{t}^{\mu})$ $=\beta(\nabla V(Z_{t}^{\mu})+\nabla W*f_{\infty}(Z_{t}^{\mu}))\cdot\nabla W*((\psi_{T-t}-1)f_{\infty})(Z_{t}^{\mu})\psi_{T-t}(Z_{t}^{\mu})\,\mathrm{d}t$ (2.73)
$\quad-\Delta W*((\psi_{T-t}-1)f_{\infty})(Z_{t}^{\mu})\psi_{T-t}(Z_{t}^{\mu})\,\mathrm{d}t+\sqrt{2\beta^{-1}}\nabla\psi_{T-t}(Z_{t}^{\mu})\cdot\mathrm{d}B_{t},$
which, integrating and taking the expectation, gives
$\operatorname{\mathbb{E}}[\psi_{T-t}(Z_{t}^{\mu})]$ $=\psi_{T}(x)-\int_{0}^{t}\operatorname{\mathbb{E}}\left[\Delta W*((\psi_{T-s}-1)f_{\infty})(Z_{s}^{\mu})\psi_{T-s}(Z_{s}^{\mu})\,\mathrm{d}s\right]$ (2.74)
$+\beta\int_{0}^{t}\operatorname{\mathbb{E}}\left[(\nabla V(Z_{s}^{\mu})+\nabla W*f_{\infty}(Z_{s}^{\mu}))\cdot\nabla W*((\psi_{T-s}-1)f_{\infty})(Z_{s}^{\mu})\psi_{T-s}(Z_{s}^{\mu})\,\mathrm{d}s\right].$
Using Corollary 2.17 we have
$\left\|(\psi_{T-s}-1)f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}\leq c_{i}e^{-a_{i}(T-s)},$ (2.75)
which implies
$\operatorname{\mathbb{E}}[\psi_{T-t}(Z_{t}^{\mu})]\leq\psi_{T}(x)+C_{i}\int_{0}^{t}e^{-a_{i}(T-s)}\operatorname{\mathbb{E}}[\psi_{T-s}(Z_{s}^{\mu})]\,\mathrm{d}s,$ (2.76)
where
$C_{i}=c_{i}\left(\left\|\Delta W\right\|_{L^{\infty}(\mathbb{T}^{d})}+\beta\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}(\left\|\nabla V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})})\right).$ (2.77)
By Grönwall’s inequality we deduce
$\operatorname{\mathbb{E}}[\psi_{T-t}(Z_{t}^{\mu})]\leq\psi_{T}(x)e^{C_{i}\int_{0}^{t}e^{-a_{i}(T-s)}\,\mathrm{d}s}\leq\psi_{T}(x)e^{C_{i}/a_{i}},$ (2.78)
which, choosing $t=T$ and by hypothesis (2.62), yields
$\psi_{T}(x)\geq\frac{1}{\kappa\Gamma}e^{-C_{i}/a_{i}}.$ (2.79)
Moreover, using again equations (2.74) and (2.75) with $t=T$, we have
$\psi_{T}(x)\leq\operatorname{\mathbb{E}}[\psi_{0}(Z_{t}^{\mu})]+C_{i}\int_{0}^{t}e^{-a_{i}(T-s)}\operatorname{\mathbb{E}}[\psi_{T-s}(Z_{s}^{\mu})]\,\mathrm{d}s,$ (2.80)
which, by equation (2.78), gives
$\psi_{T}(x)\leq\operatorname{\mathbb{E}}[\psi_{0}(Z_{t}^{\mu})]+C_{i}\psi_{T}(x)e^{C_{i}/a_{i}}\int_{0}^{t}e^{-a_{i}(T-s)}\,\mathrm{d}s\leq\kappa\Gamma+\frac{C_{i}}{a_{i}}\psi_{T}(x)e^{C_{i}/a_{i}}.$ (2.81)
Combining bounds (2.79) and (2.81), we obtain for all $T>0$
$\frac{1}{\kappa\Gamma e^{C_{i}/a_{i}}}\leq\psi_{T}(x)\leq\frac{\kappa\Gamma}{1-(C_{i}/a_{i})e^{C_{i}/a_{i}}},$ (2.82)
where we remark that, due to the assumptions on $a_{i}$ and $C_{i}$, the inequalities make sense since the left-hand side and the right-hand side are always smaller and greater than one, respectively. Notice also that it holds
$\kappa\Gamma e^{C_{i}/a_{i}}\leq\frac{\kappa\Gamma}{1-(C_{i}/a_{i})e^{C_{i}/a_{i}}}.$ (2.83)
Therefore, the desired results follow from the bounds (2.71) and (2.82) by applying the Holley–Stroock perturbation lemma (see, e.g., [33, Lemma 2.3]) and Lemma 2.12. ∎
Using the LSI, we can finally prove that the measures $\mu_{t}$ and $\nu_{t}$ are exponentially close in time, in relative entropy. The argument is similar to the proof of Theorem 2.7.
Theorem 2.19.
Let the assumptions of Lemma 2.18 be satisfied. Then, it holds for $i=1,2$
$\mathcal{H}(\mu_{t}|\nu_{t})\leq\sigma_{\Xi}(t)\coloneqq\begin{cases}\left(\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta^{2}c_{i}\Xi\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}^{2}}{2(2-\beta a_{i}\Xi)}\right)e^{-a_{i}t},&\text{if }\Xi<\frac{2}{a_{i}\beta},\\ \left(\mathcal{H}(\mu_{0}|\nu_{0})+\frac{1}{2}\beta c_{i}\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}^{2}t\right)e^{-\frac{2}{\beta\Xi}t},&\text{if }\Xi=\frac{2}{a_{i}\beta},\\ \left(\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta^{2}c_{i}\Xi\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}^{2}}{2(\beta a_{i}\Xi-2)}\right)e^{-\frac{2}{\beta\Xi}t},&\text{if }\Xi>\frac{2}{a_{i}\beta},\end{cases}$ (2.84)
where $a_{i}$ and $c_{i}$ are defined in Corollary 2.17.
Proof.
By [39, Lemma 3.1] we have the estimate
$\mathcal{H}(\mu_{t}|\nu_{t})+\frac{1}{2\beta}\int_{0}^{t}\mathcal{I}(\mu_{s}|\nu_{s})\,\mathrm{d}s\leq\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta}{2}\int_{0}^{t}\int_{\mathbb{T}^{d}}\left\lvert(\nabla W*(f_{s}-f_{\infty}))(x)\right\rvert^{2}f_{s}(x)\,\mathrm{d}x\,\mathrm{d}s,$ (2.85)
which, due to Lemma 2.18, implies
$\mathcal{H}(\mu_{t}|\nu_{t})+\frac{2}{\beta\Xi}\int_{0}^{t}\mathcal{H}(\mu_{s}|\nu_{s})\,\mathrm{d}s\leq\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta}{2}\int_{0}^{t}\int_{\mathbb{T}^{d}}\left\lvert(\nabla W*(f_{s}-f_{\infty}))(x)\right\rvert^{2}f_{s}(x)\,\mathrm{d}x\,\mathrm{d}s.$ (2.86)
Then, Corollary 2.17 gives
$\left\lvert(\nabla W*(f_{s}-f_{\infty}))(x)\right\rvert^{2}\leq\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}^{2}\left\|f_{s}-f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}^{2}\leq c_{i}\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}^{2}e^{-a_{i}s},$ (2.87)
which yields
$\mathcal{H}(\mu_{t}|\nu_{t})\leq\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta c_{i}\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}^{2}}{2}\int_{0}^{t}e^{-a_{i}s}\,\mathrm{d}s-\frac{2}{\beta\Xi}\int_{0}^{t}\mathcal{H}(\mu_{s}|\nu_{s})\,\mathrm{d}s.$ (2.88)
Applying [39, Lemma A.1], we have
$\mathcal{H}(\mu_{t}|\nu_{t})$ $\leq\mathcal{H}(\mu_{0}|\nu_{0})e^{-\frac{2}{\beta\Xi}t}+\frac{\beta c_{i}\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}^{2}}{2}\int_{0}^{t}e^{-\frac{2}{\beta\Xi}(t-s)}e^{-a_{i}s}\,\mathrm{d}s,$ (2.89)
$=\mathcal{H}(\mu_{0}|\nu_{0})e^{-\frac{2}{\beta\Xi}t}+\frac{\beta c_{i}\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}^{2}}{2}e^{-\frac{2}{\beta\Xi}t}\int_{0}^{t}e^{\frac{2-a_{i}\beta\Xi}{\beta\Xi}s}\,\mathrm{d}s.$
Therefore, if $\Xi=2/(a_{i}\beta)$, we deduce
$\mathcal{H}(\mu_{t}|\nu_{t})\leq\left(\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta c_{i}\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}^{2}}{2}t\right)e^{-\frac{2}{\beta\Xi}t},$ (2.90)
otherwise we get
$\mathcal{H}(\mu_{t}|\nu_{t})\leq\mathcal{H}(\mu_{0}|\nu_{0})e^{-\frac{2}{\beta\Xi}t}+\frac{\beta^{2}c_{i}\Xi\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}^{2}}{2(2-\beta a_{i}\Xi)}\left(e^{-a_{i}t}-e^{-\frac{2}{\beta\Xi}t}\right),$ (2.91)
which concludes the proof. ∎
Similarly to the case where the state space is the whole space, the following convergence in Wasserstein distance is a direct consequence of the Talagrand’s transportation inequality.
Corollary 2.20.
Let the assumptions of Theorem 2.19 be satisfied. Then, it holds
$\mathcal{W}_{2}(\mu_{t},\nu_{t})\leq\sqrt{\Xi\sigma_{\Xi}(t)},$ (2.92)
where $\mathcal{W}_{2}$ denotes the 2-Wasserstein distance and $\sigma_{\Xi}(t)$ is defined in Theorem 2.19.
Proof.
The desired result follows from Theorem 2.19 using [48, Theorem 1]. ∎
3 Applications
In this section, we present two applications in which the linearization procedure can be used. From the analysis presented in the previous section, the law of the solution of the linearized McKean SDE $\nu_{t}$ is close to the law of the nonlinear process $\mu_{t}$, and their distance vanishes exponentially fast as $t\to\infty$. Therefore, the linear process $Y_{t}$ can provide a reliable approximation when studying properties that depend on the long-time behaviour of the nonlinear dynamics $X_{t}$. We first study the maximum likelihood estimator for the mean field dynamics. Then, we compute the diffusion coefficient of the limiting Brownian motion in the diffusive-mean field limit of weakly interacting diffusions in periodic potentials.
3.1 Maximum likelihood estimation
We consider the case where the confining and interaction potentials $V$ and $W$ depend on unknown parameters. Our goal is to infer these parameters by observing a trajectory of the mean field SDE (1.2). Hence, consider the equation on the whole space $\mathbb{R}^{d}$ and in the time horizon $[0,T]$
$\mathrm{d}X_{t}=-\nabla V(X_{t};\theta)\,\mathrm{d}t-(\nabla W(\cdot;\theta)*f_{t}(\cdot;\theta))(X_{t})\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},$ (3.1)
where $\theta\in\Theta$ is the vector of unknown parameters and $\Theta$ is the admissible set. The noise strength is assumed to be known. We denote by $\theta_{0}$ the exact value and follow a maximum likelihood approach [34, 7, 24]. In particular, by the Girsanov theorem, the log-likelihood is
$\mathfrak{L}_{T}(\theta)$ $=-\frac{\beta}{2}\int_{0}^{T}\left(\nabla V(X_{t};\theta)+(\nabla W(\cdot;\theta)*f_{t}(\cdot;\theta))(X_{t})\right)^{\top}\,\mathrm{d}X_{t}$ (3.2)
$\quad-\frac{\beta}{4}\int_{0}^{T}\left\lvert\nabla V(X_{t};\theta)+(\nabla W(\cdot;\theta)*f_{t}(\cdot;\theta))(X_{t})\right\rvert^{2}\,\mathrm{d}t,$
and the MLE is defined as
$\widehat{\theta}_{T}=\operatorname{argmax}_{\theta\in\Theta}\mathfrak{L}_{T}(\theta).$ (3.3)
We will use the following notation.
$b_{\theta}(t,x)$ $\coloneqq-\nabla V(x;\theta)-(\nabla W(\cdot;\theta)*f_{t}(\cdot;\theta))(x),$ (3.4)
$\widetilde{b}_{\theta}(x)$ $\coloneqq-\nabla V(x;\theta)-(\nabla W(\cdot;\theta)*f_{\infty}(\cdot;\theta))(x),$
and notice that maximizing $\mathfrak{L}_{T}$ is equivalent to maximize the quantity
$\ell_{T}(\theta)=\frac{\beta}{2}\int_{0}^{T}\left(b_{\theta}(t,X_{t})-b_{\theta_{0}}(t,X_{t})\right)^{\top}\,\mathrm{d}B_{t}-\frac{\beta}{4}\int_{0}^{T}\left\lvert b_{\theta}(t,X_{t})-b_{\theta_{0}}(t,X_{t})\right\rvert^{2}\,\mathrm{d}t.$ (3.5)
In particular, the MLE is defined as
$\widehat{\theta}_{T}=\operatorname{argmax}_{\theta\in\Theta}\mathfrak{L}_{T}(\theta)=\operatorname{argmax}_{\theta\in\Theta}\ell_{T}(\theta).$ (3.6)
We now define the linearized version of the log-likelihood around the invariant measure $f_{\infty}$,
$\widetilde{\mathfrak{L}}_{T}(\theta)$ $=-\frac{\beta}{2}\int_{0}^{T}\left(\nabla V(X_{t};\theta)+(\nabla W(\cdot;\theta)*f_{\infty}(\cdot;\theta))(X_{t})\right)^{\top}\,\mathrm{d}X_{t}$ (3.7)
$\quad-\frac{\beta}{4}\int_{0}^{T}\left\lvert\nabla V(X_{t};\theta)+(\nabla W(\cdot;\theta)*f_{\infty}(\cdot;\theta))(X_{t})\right\rvert^{2}\,\mathrm{d}t,$
and the analogue of (3.5)
~$\widetilde{\ell}_{T}(\theta)\coloneqq\frac{\beta}{2}\int_{0}^{T}\left(\widetilde{b}_{\theta}(X_{t})-b_{\theta_{0}}(t,X_{t})\right)^{\top}\,\mathrm{d}B_{t}-\frac{\beta}{4}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta}(X_{t})-b_{\theta_{0}}(t,X_{t})\right\rvert^{2}\,\mathrm{d}t.$ (3.8)
The linearized MLE is therefore given by
$\widetilde{\theta}_{T}=\operatorname{argmax}_{\theta\in\Theta}\widetilde{\mathfrak{L}}_{T}(\theta)=\operatorname{argmax}_{\theta\in\Theta}\widetilde{\ell}_{T}(\theta).$ (3.9)
It is already known that the MLE $\widehat{\theta}_{T}$, which is computed on a path from the solution to the McKean SDE, is asymptotically unbiased [40, Theorem 3.1], meaning that
[…]
and we aim to show that the same result holds true for the linearized MLE $\widetilde{\theta}_{T}$.
3.1.1 Asymptotic unbiasedness of the linearized MLE
In this section, we show that replacing the density of the process by its invariant measure in the calculation of the likelihood function, while observing a path of the McKean SDE, does not affect the resulting MLE. We emphasize the fact that we keep the same observation model, i.e., that the “linearized” likelihood is still evaluated at a path of the nonlinear process. We will make the following assumptions.
Assumption 3.1.
The following properties hold true.
  1. (i)
    The admissible set $\Theta$ is compact.
  2. (ii)
    The gradients of the confining and interaction potentials are Lipschitz in $\theta$, i.e., there exist $L_{V},L_{W}\colon\mathbb{R}^{d}\to\mathbb{R}$, $L_{V},L_{W}>0$, such that for all $\theta_{1},\theta_{2}\in\Theta$ and for all $x\in\mathbb{R}^{d}$
    $\left\lvert\nabla V(x;\theta_{1})-\nabla V(x;\theta_{2})\right\rvert$ $\leq L_{V}(x)\left\lvert\theta_{1}-\theta_{2}\right\rvert,$ (3.11)
    $\left\lvert\nabla W(x;\theta_{1})-\nabla W(x;\theta_{2})\right\rvert$ $\leq L_{W}(x)\left\lvert\theta_{1}-\theta_{2}\right\rvert,$
    where $L_{V}(x)$ and $L_{W}(x)$ are polynomially bounded.
  3. (iii)
    The function
    $\theta\mapsto\int_{\mathbb{R}^{d}}\left\lvert\widetilde{b}_{\theta}(x)-\widetilde{b}_{\theta_{0}}(x)\right\rvert^{2}f_{\infty}(x;\theta)\,\mathrm{d}x$ (3.12)
    admits only one zero at $\theta=\theta_{0}$.
Remark 3.2.
The condition in Assumption 3.1(iii) is necessary to guarantee the identifiability of the unknown parameter $\theta_{0}$, and it is related to the nondegeneracy of the Fisher information matrix in [24].
The main result of this section is the asymptotic unbiasedness of $\widetilde{\theta}_{T}$, when evaluated at a path of the McKean SDE.
Theorem 3.3.
Under Assumptions 2.3 and 3.1, it holds
[…]
Remark 3.4.
We reiterate that our observation model is that of a single (nonstationary) path of the McKean SDE. Ideally, we would like to consider the case where we observe one or more paths of the interacting particle system. To show asymptotic unbiasedness of the linearized MLE in this setting, we would need to approximate the path of the nonlinear process by a path of the particle system, using the uniform propagation of chaos property. We will omit the details. Needless to say, in the numerical experiments we observe a single trajectory of the particle system.
Our theoretical analysis is inspired by the proof of [11, Corollary 4.1], where the linear case is treated. In fact, in our approach the likelihood function is computed from the linear process, and the nonlinearity only appears in the observation model. The proof of Theorem 3.3 relies on the following preliminary results.
Lemma 3.5.
Under Assumptions 2.3 and 3.1, it holds for all $\theta\in\Theta$
[…]
Proof.
Let us denote by $I_{T}$ the quantity in the left-hand side of the result, and notice that the following decomposition holds true
$I_{T}$ $=\frac{1}{T}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta}(X_{t})-b_{\theta_{0}}(t,X_{t})\right\rvert^{2}\,\mathrm{d}t$ (3.15)
$=\frac{1}{T}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta}(X_{t})-\widetilde{b}_{\theta_{0}}(X_{t})\right\rvert^{2}\,\mathrm{d}t$
$\quad+\frac{1}{T}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta_{0}}(X_{t})-b_{\theta_{0}}(t,X_{t})\right\rvert^{2}\,\mathrm{d}t$
$\quad+\frac{2}{T}\int_{0}^{T}\left(\widetilde{b}_{\theta}(X_{t})-\widetilde{b}_{\theta_{0}}(X_{t})\right)^{\top}\left(\widetilde{b}_{\theta_{0}}(X_{t})-b_{\theta_{0}}(t,X_{t})\right)\,\mathrm{d}t$
$=I_{T}^{(1)}+I_{T}^{(2)}+I_{T}^{(3)}.$
By the ergodic theorem we first have
[…]
We then consider $I_{T}^{(2)}$ and, due to the definition (3.4) and the Cauchy–Schwarz inequality, we get
$I_{T}^{(2)}$ $\leq\frac{1}{T}\int_{0}^{T}\left(\int_{\mathbb{R}^{d}}\left\lvert\nabla W(X_{t}-y;\theta_{0})\right\rvert\left\lvert f_{t}(y;\theta_{0})-f_{\infty}(y;\theta_{0})\right\rvert\,\mathrm{d}y\right)^{2}\,\mathrm{d}t$ (3.17)
$\leq\frac{1}{T}\int_{0}^{T}\left(\int_{\mathbb{R}^{d}}\left\lvert\nabla W(X_{t}-y;\theta_{0})\right\rvert^{2}\left(f_{t}(y;\theta_{0})+f_{\infty}(y;\theta_{0})\right)\,\mathrm{d}y\right)\,\mathrm{d}t$
×$\hskip 28.45274pt\times\left\|f_{t}(\cdot;\theta_{0})-f_{\infty}(\cdot;\theta_{0})\right\|_{L^{1}(\mathbb{R}^{d})},$
which, by equation (2.8), implies
$I_{T}^{(2)}\leq\frac{K}{T}e^{-\frac{1}{2}\alpha t}\int_{0}^{T}\left(\int_{\mathbb{R}^{d}}\left\lvert\nabla W(X_{t}-y;\theta_{0})\right\rvert^{2}\left(f_{t}(y;\theta_{0})+f_{\infty}(y;\theta_{0})\right)\,\mathrm{d}y\right)\,\mathrm{d}t.$ (3.18)
Then, by the fact that $\nabla W$ is polynomially bounded, the uniform boundedness of the moments, and the ergodic theorem, we deduce that $I_{T}^{(2)}\to 0$ almost surely. Moreover, by Cauchy–Schwarz inequality we have
$I_{T}^{(3)}\leq 2\sqrt{I_{T}^{(1)}I_{T}^{(2)}},$ (3.19)
which yields that $I_{T}^{(3)}$ converges to zero almost surely. The desired result then follows from the convergence of the terms $I_{T}^{(1)},I_{T}^{(2)},I_{T}^{(3)}$. ∎
Corollary 3.6.
Under Assumptions 2.3 and 3.1, it holds
[…]
Proof.
The desired result follows from Lemma 3.5 by replacing $\theta$ with $\theta_{0}$. ∎
Lemma 3.7.
Under Assumptions 2.3 and 3.1, it holds
[…]
Proof.
Let us define the function
$e(T,\theta)=\frac{1}{T}\int_{0}^{T}\left(\widetilde{b}_{\theta}(X_{t})-b_{\theta_{0}}(t,X_{t})\right)\,\mathrm{d}B_{t},$ (3.22)
which is uniformly continuous in $\theta$ on $\Theta$, uniformly in $T\in[1,\infty)$, with probability $1$, by [11, Theorem 3.1]. Then, by Lemma 3.5 and the strong law of large numbers for martingales we deduce that for all $\theta\in\Theta$
[…]
Finally, the result is obtained following the same argument as in the proof of [11, Lemma 4.1]. ∎
Lemma 3.8.
Under Assumptions 2.3 and 3.1, there exists a constant $L>0$, independent of $T$, such that for all $\theta_{1},\theta_{2}\in\Theta$
$\left\lvert\frac{1}{T}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta_{1}}(X_{t})-b_{\theta_{0}}(t,X_{t})\right\rvert^{2}\,\mathrm{d}t-\frac{1}{T}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta_{2}}(X_{t})-b_{\theta_{0}}(t,X_{t})\right\rvert^{2}\,\mathrm{d}t\right\rvert\leq L\left\lvert\theta_{1}-\theta_{2}\right\rvert.$ (3.24)
Proof.
In order to show the result, we define the function
$h(T,\theta)=\frac{1}{T}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta}(X_{t})-b_{\theta_{0}}(t,X_{t})\right\rvert^{2}\,\mathrm{d}t,$ (3.25)
and show that it is Lipschitz in $\theta$, uniformly in $T$. By the Cauchy–Schwarz inequality, we have for a constant $B>0$
$\left\lvert h(T,\theta_{1})-h(T,\theta_{2})\right\rvert=\left\lvert\frac{1}{T}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta_{1}}(X_{t})-b_{\theta_{0}}(t,X_{t})\right\rvert^{2}\,\mathrm{d}t-\frac{1}{T}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta_{2}}(X_{t})-b_{\theta_{0}}(t,X_{t})\right\rvert^{2}\,\mathrm{d}t\right\rvert$ (3.26)
$\hskip 42.67912pt\leq\frac{1}{T}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta_{1}}(X_{t})-\widetilde{b}_{\theta_{2}}(X_{t})\right\rvert\left\lvert\widetilde{b}_{\theta_{1}}(X_{t})+\widetilde{b}_{\theta_{2}}(X_{t})-2b_{\theta_{0}}(t,X_{t})\right\rvert\,\mathrm{d}t$
$\hskip 42.67912pt\leq\left(\frac{1}{T}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta_{1}}(X_{t})-\widetilde{b}_{\theta_{2}}(X_{t})\right\rvert^{2}\right)^{1/2}\left(\frac{1}{T}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta_{1}}(X_{t})+\widetilde{b}_{\theta_{2}}(X_{t})-2b_{\theta_{0}}(t,X_{t})\right\rvert^{2}\right)^{1/2}$
$\hskip 42.67912pt\leq B\left(\frac{1}{T}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta_{1}}(X_{t})-\widetilde{b}_{\theta_{2}}(X_{t})\right\rvert^{2}\right)^{1/2},$
where the uniform bound in time of the last inequality is due to Lemma 3.5. By the definition of $\widetilde{b}_{\theta}$ in equation (3.4) we obtain
$\left\lvert h(T,\theta_{1})-h(T,\theta_{2})\right\rvert^{2}$ $\leq\frac{B}{T}\int_{0}^{T}\left\lvert\nabla V(X_{t};\theta_{1})-\nabla V(X_{t};\theta_{2})\right\rvert^{2}\,\mathrm{d}t$ (3.27)
$+\frac{B}{T}\int_{0}^{T}\left(\int_{\mathbb{R}^{d}}\left\lvert\nabla W(X_{t}-y;\theta_{1})\right\rvert\left\lvert f_{\infty}(y;\theta_{1})-f_{\infty}(y;\theta_{2})\right\rvert\,\mathrm{d}y\right)^{2}\,\mathrm{d}t$
$+\frac{B}{T}\int_{0}^{T}\left(\int_{\mathbb{R}^{d}}\left\lvert\nabla W(X_{t}-y;\theta_{1})-\nabla W(X_{t}-y;\theta_{2})\right\rvert f_{\infty}(y;\theta_{2})\,\mathrm{d}y\right)^{2}\,\mathrm{d}t,$
which by Cauchy–Schwarz inequality gives
$\left\lvert h(T,\theta_{1})-h(T,\theta_{2})\right\rvert^{2}$ $\leq\frac{B}{T}\int_{0}^{T}\left\lvert\nabla V(X_{t};\theta_{1})-\nabla V(X_{t};\theta_{2})\right\rvert^{2}\,\mathrm{d}t$ (3.28)
$+\frac{B}{T}\int_{0}^{T}\left(\int_{\mathbb{R}^{d}}\left\lvert\nabla W(X_{t}-y;\theta_{1})\right\rvert^{2}(f_{\infty}(y;\theta_{1})+f_{\infty}(y;\theta_{2}))\,\mathrm{d}y\right)$
×$\hskip 31.2982pt\times\left(\int_{\mathbb{R}^{d}}\left\lvert f_{\infty}(y;\theta_{1})-f_{\infty}(y;\theta_{2})\right\rvert\,\mathrm{d}y\right)\,\mathrm{d}t$
$+\frac{B}{T}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left\lvert\nabla W(X_{t}-y;\theta_{1})-\nabla W(X_{t}-y;\theta_{2})\right\rvert^{2}f_{\infty}(y;\theta_{2})\,\mathrm{d}y\,\mathrm{d}t.$
Moreover, since $\nabla V$ and $\nabla W$ are Lipschitz in $\theta$ by Assumption 3.1(ii) and polynomially bounded in $x$, using the uniform boundedness of the moments and the ergodic theorem, and due to the differentiability of the invariant measure with respect to $\theta$ [9], we have that there exists a constant $L>0$, independent of $T$, such that
$\left\lvert h(T,\theta_{1})-h(T,\theta_{2})\right\rvert^{2}\leq L^{2}\left\lvert\theta_{1}-\theta_{2}\right\rvert^{2},$ (3.29)
which yields the desired result. ∎
We can finally prove Theorem 3.3.
Proof of Theorem 3.3.
By definition, the estimator $\widetilde{\theta}_{T}$ satisfies for all $\theta\in\Theta$
$\frac{1}{T}\widetilde{\ell}_{T}(\widetilde{\theta}_{T})\geq\frac{1}{T}\widetilde{\ell}_{T}(\theta),$ (3.30)
and, in particular, taking $\theta=\theta_{0}$, we get by equation (3.8)
~$\frac{\widetilde{\ell}_{T}(\widetilde{\theta}_{T})}{T}\geq\frac{\widetilde{\ell}_{T}(\theta_{0})}{T}=\frac{\beta}{2T}\int_{0}^{T}\left(\widetilde{b}_{\theta_{0}}(X_{t})-b_{\theta_{0}}(t,X_{t})\right)^{\top}\,\mathrm{d}B_{t}-\frac{\beta}{4T}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta_{0}}(X_{t})-b_{\theta_{0}}(t,X_{t})\right\rvert^{2}\,\mathrm{d}t,$ (3.31)
which implies
$\frac{\beta}{4T}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta_{0}}(X_{t})-b_{\theta_{0}}(t,X_{t})\right\rvert^{2}\,\mathrm{d}t+\left[\frac{\beta}{2T}\int_{0}^{T}\left(\widetilde{b}_{\theta}(X_{t})-b_{\theta_{0}}(t,X_{t})\right)^{\top}\,\mathrm{d}B_{t}\right]_{\theta=\widetilde{\theta}_{T}}$ (3.32)
$\hskip 42.67912pt-\frac{\beta}{2T}\int_{0}^{T}\left(\widetilde{b}_{\theta_{0}}(X_{t})-b_{\theta_{0}}(t,X_{t})\right)^{\top}\,\mathrm{d}B_{t}\geq\frac{\beta}{4T}\int_{0}^{T}\left\lvert\widetilde{b}_{\widetilde{\theta}_{T}}(X_{t})-b_{\theta_{0}}(t,X_{t})\right\rvert^{2}\,\mathrm{d}t\geq 0.$
Applying Corollaries 3.6 and 3.7 we obtain
[…]
Let us now define the function $h=h(T,\theta)$ as in equation (3.25), which is Lipschitz in $\theta$, uniformly in $T$, by Lemma 3.8, and notice that
[…]
Therefore, we deduce that $\widetilde{\theta}_{T}\to\Omega$ as $T\to\infty$ almost surely, where
$\Omega=\left\{\theta\in\Theta\colon\lim_{T\to\infty}h(T,\theta)=0\right\},$ (3.35)
which by definition (3.25) and Lemma 3.5 coincides with
$\Omega=\left\{\theta\in\Theta\colon\int_{\mathbb{R}^{d}}\left\lvert\widetilde{b}_{\theta}(x)-\widetilde{b}_{\theta_{0}}(x)\right\rvert^{2}f_{\infty}(x)\,\mathrm{d}x=0\right\}.$ (3.36)
Finally, by the identifiability condition in Assumption 3.1(iii) we get that $\Omega=\{\theta_{0}\}$, and the desired result follows. ∎
3.1.2 Numerical examples
Figure 1: Comparison between the MLE $\widehat{\theta}_{T}$ and the linearized MLE $\widetilde{\theta}_{T}$ for different values of the final time of observation $T\in[0,10^{3}]$, considering the quadratic interaction potential and two different examples of confining potentials.
We now implement the linearized MLE and test it on two numerical examples. In all the experiments presented in this section we consider deterministic initial conditions, $\mu_{0}=\nu_{0}=\delta_{1}$. For the discretization we use the Euler–Maruyama method with a time step $\Delta t=10^{-3}$. We will also set $\beta=1$.
We first consider the one-dimensional mean field Ornstein–Uhlenbeck process, i.e., we set $V(x;\theta)=\theta x^{2}/2$ and $W(x;\theta)=x^{2}/2$ in equation (3.1), and therefore the McKean SDE reads
$\mathrm{d}X_{t}=-\theta X_{t}\,\mathrm{d}t-(X_{t}-\operatorname{\mathbb{E}}[X_{t}])\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},$ (3.37)
where we set the exact value of the unknown parameter to be $\theta_{0}=1$. We take the expectation in (3.37) to obtain
$\operatorname{\mathbb{E}}[X_{t}]=\operatorname{\mathbb{E}}[X_{0}]-\theta\int_{0}^{t}\operatorname{\mathbb{E}}[X_{s}]\,\mathrm{d}s,$ (3.38)
which implies
$\operatorname{\mathbb{E}}[X_{t}]=\operatorname{\mathbb{E}}[X_{0}]e^{-\theta t}.$ (3.39)
Therefore, the McKean SDE (3.37) can be rewritten as
$\mathrm{d}X_{t}=-\theta X_{t}\,\mathrm{d}t-(X_{t}-\operatorname{\mathbb{E}}[X_{0}]e^{-\theta t})\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},$ (3.40)
and, since its invariant measure is $\mu_{\infty}=\mathcal{N}(0,1/(\beta(\theta+1)))$, the linearized process $Y_{t}$ satisfies
$\mathrm{d}Y_{t}=-(\theta+1)Y_{t}\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t}.$ (3.41)
Minimizing the log-likelihood $\mathfrak{L}_{T}$ and the linearized log-likelihood $\widetilde{\mathfrak{L}}_{T}$ in equations (3.2) and (3.7), respectively, we deduce that the MLE $\widehat{\theta}_{T}$ solves the nonlinear equation
00 0 $=-\int_{0}^{T}X_{t}\,\mathrm{d}X_{t}-(\widehat{\theta}_{T}+1)\int_{0}^{T}X_{t}^{2}\,\mathrm{d}t-\operatorname{\mathbb{E}}[X_{0}]\int_{0}^{T}te^{-\widehat{\theta}_{T}t}\,\mathrm{d}X_{t}$ (3.42)
$\quad+(\widehat{\theta}_{T}+1)\operatorname{\mathbb{E}}[X_{0}]\int_{0}^{T}te^{-\widehat{\theta}_{T}t}X_{t}\,\mathrm{d}t+\operatorname{\mathbb{E}}[X_{0}]\int_{0}^{T}e^{-\widehat{\theta}_{T}t}X_{t}\,\mathrm{d}t-\operatorname{\mathbb{E}}[X_{0}]^{2}\int_{0}^{T}te^{-2\widehat{\theta}_{T}t}\,\mathrm{d}t,$
while the linearized MLE $\widetilde{\theta}_{T}$ is given by the closed form expression
$\widetilde{\theta}_{T}=-\frac{\int_{0}^{T}X_{t}\,\mathrm{d}X_{t}}{\int_{0}^{T}X_{t}^{2}\,\mathrm{d}t}-1.$ (3.43)
In the first plot in Figure 1, we compare the two estimators for different values of the final time of observation $T\in[0,10^{3}]$, and notice that, except for a short time interval, the difference between $\widehat{\theta}_{T}$ and $\widetilde{\theta}_{T}$ is negligible. In particular, they both provide an accurate approximation of the exact unknown.
In the second test case, we consider the bistable confining potential below the phase transition, i.e., when we have uniqueness of the invariant measure; see [19, 51] for details. Let $W(x;\theta)=x^{2}/2$ and $V(x;\theta)=\theta(x^{4}/4-x^{2}/2)$ with true parameter $\theta_{0}=1$. Then, the invariant measure $\mu_{\infty}$ has density
$f_{\infty}(x)=\frac{e^{-\theta\frac{x^{4}}{4}-(1-\theta)\frac{x^{2}}{2}}}{\int_{\mathbb{R}}e^{-\theta\frac{z^{4}}{4}-(1-\theta)\frac{z^{2}}{2}}\,\mathrm{d}z}.$ (3.44)
The nonlinear and linear, respectively, processes $X_{t}$ and $Y_{t}$ solve the SDEs
$\mathrm{d}X_{t}$ $=-\theta(X_{t}^{3}-X_{t})\,\mathrm{d}t-(X_{t}-\operatorname{\mathbb{E}}[X_{t}])\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},$ (3.45)
$\mathrm{d}Y_{t}$ $=-\theta X_{t}^{3}\,\mathrm{d}t-(1-\theta)X_{t}\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t}.$
Assuming to know, in addition to the trajectory $(X_{t})_{t\in[0,T]}$, also its expectation at all times $(\operatorname{\mathbb{E}}[X_{t}])_{t\in[0,T]}$, we can write the MLE as
$\widehat{\theta}_{T}=-\frac{\int_{0}^{T}(X_{t}^{3}-X_{t})\,\mathrm{d}X_{t}+\int_{0}^{T}(X_{t}-\operatorname{\mathbb{E}}[X_{t}])(X_{t}^{3}-X_{t})\,\mathrm{d}t}{\int_{0}^{T}(X_{t}^{3}-X_{t})^{2}\,\mathrm{d}t}.$ (3.46)
On the other hand, the linearized MLE is given by the simpler expression
$\widetilde{\theta}_{T}=-\frac{\int_{0}^{T}(X_{t}^{3}-X_{t})\,\mathrm{d}X_{t}+\int_{0}^{T}X_{t}(X_{t}^{3}-X_{t})\,\mathrm{d}t}{\int_{0}^{T}(X_{t}^{3}-X_{t})^{2}\,\mathrm{d}t},$ (3.47)
which does not need the expectation $\operatorname{\mathbb{E}}[X_{t}]$. We note that, in this case, McKean’s SDE must be approximated through the interacting particle system (1.1). In particular, we use $N=500$ particles, and a path of the mean field SDE is approximated by a path of one of the trajectories of the particle system. Due to exchangeability, it does not matter which particle we choose. We notice that assuming to know the expectation $\operatorname{\mathbb{E}}[X_{t}]$ at all times implies that all the interacting particle system must be observed. On the contrary, for the linearized MLE, only the trajectory of a single particle is required. The behavior of the two estimators for different values of the final time of observation $T\in[0,10^{3}]$ is shown in the second plot in Figure 1. As expected, the two estimators provide us with almost identical values for the estimated parameter, after a short initial transient time.
3.2 The diffusive-mean field limit
We revisit the problem of studying the joint diffusive-mean field limit for weakly interacting diffusions in periodic confining and interaction potentials that was studied in [22]. As discussed in that paper, the two limits of sending $\varepsilon\to 0$ and $N\to\infty$ commute only when the mean field dynamics on the torus does not exhibit phase transitions. Here, we study only the “easy” order of taking the two limits, first $N\to\infty$ and then sending $\varepsilon$ to 00, and only under the assumption that the mean field SDE has a unique steady state. A full analysis of the problem based on the linearized SDE, for both the overdamped and the underdamped dynamics, will be presented elsewhere.
We recall the McKean SDE (1.2)
$\mathrm{d}X_{t}=-\nabla V(X_{t})\,\mathrm{d}t-(\nabla W*f_{t})(X_{t})\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},$ (3.48)
where now the state space is $\mathbb{R}^{d}$, but the confining and interaction potentials $V$ and $W$ are assumed to be periodic with period the $d$-dimensional torus $\mathbb{T}^{d}$. First, notice that we have
$(\nabla W*f_{t})(x)=\int_{\mathbb{R}^{d}}\nabla W(x-y)f_{t}(y)\,\mathrm{d}y=\sum_{k\in\mathbb{Z}^{d}}\int_{k+\mathbb{T}^{d}}\nabla W(x-y)f_{t}(y)\,\mathrm{d}y,$ (3.49)
which, by the change of variables $z=y-k$ and the periodicity of $W$, implies
$(\nabla W*f_{t})(x)=\sum_{k\in\mathbb{Z}^{d}}\int_{\mathbb{T}^{d}}\nabla W(x-z-k)f_{t}(z+k)\,\mathrm{d}z=\int_{\mathbb{T}^{d}}\nabla W(x-z)\sum_{k\in\mathbb{Z}^{d}}f_{t}(z+k)\,\mathrm{d}z.$ (3.50)
Then, by defining the density $\phi_{t}\colon\mathbb{T}^{d}\to\mathbb{R}$ as
$\phi_{t}(x)=\sum_{k\in\mathbb{Z}^{d}}f_{t}(x+k),$ (3.51)
we get
$(\nabla W*f_{t})(x)=\int_{\mathbb{T}^{d}}\nabla W(x-z)\phi_{t}(z)\,\mathrm{d}z=(\nabla W*\phi_{t})(x).$ (3.52)
Moreover, considering the wrapping of the process $X_{t}$ on the torus, it is possible to prove that $\phi_{t}$ converges to the solution $\phi_{\infty}\colon\mathbb{T}^{d}\to\mathbb{R}$ of the stationary Fokker–Planck equation on $\mathbb{T}^{d}$
$\nabla\cdot\left(\left(\nabla V+\nabla W*\phi_{\infty}\right)\phi_{\infty}\right)+\beta^{-1}\Delta\phi_{\infty}=0.$ (3.53)
This nonlinear and nonlocal elliptic PDE can be reformulated as the integral equation (1.4); see [60, 25]. In [22], letting $\varepsilon>0$, the diffusive rescaling $\widetilde{X}_{t}^{\varepsilon}\coloneqq\varepsilon X_{t/\varepsilon^{2}}$ is considered and its limit as $\varepsilon$ vanishes is studied. In particular, in [22, Theorem 1.10], under the assumption that the mean field dynamics on the torus is geometrically ergodic, the following functional central limit theorem (invariance principle) is proved. It holds
$\lim_{\varepsilon\to 0}\widetilde{X}_{t}^{\varepsilon}=\sqrt{2\beta^{-1}\mathcal{D}}B_{t},\qquad\text{in law},$ (3.54)
where $B_{t}$ is a standard $d$-dimensional Brownian motion and the matrix $\mathcal{D}$ is given by the standard homogenization formula [49, Chapter 13]
$\mathcal{D}=\operatorname{\mathbb{E}}^{\phi_{\infty}}[(I+\nabla\Phi(X))(I+\nabla\Phi(X))^{\top}],$ (3.55)
where the superscript denotes the fact that the expectation is computed with respect to the invariant measure with density $\phi_{\infty}$. The function $\Phi\colon\mathbb{T}^{d}\to\mathbb{R}^{d}$ in the equation above is the unique mean zero solution of the Poisson problem on $\mathbb{T}^{d}$
$\nabla\cdot\left(\phi_{\infty}\nabla\Phi\right)=\nabla\phi_{\infty},$ (3.56)
with periodic boundary conditions. The proof of this result is based on the standard methodology: application of Itô’s formula to the solution of the (nolinear) Poisson equation, decomposition of the rescaled process into a martingale part and a reminder, use of PDE estimates on the Poisson equation together with the martingale central limit theorem. We remark that, for the proof of the invariance principle in our setting, it is sufficient to prove a central limit theorem of the form
$\lim_{t\to\infty}\frac{1}{\sqrt{t}}X_{t}=\mathcal{N}(0,2\beta^{-1}\mathcal{D}),\qquad\text{in law},$ (3.57)
as done in [22, Lemma 2.3]. The invariance principle then follows from standard arguments; see, e.g., [47, 37]. We also remark that equation (3.57) is also implied by equation (3.54), by setting $t=1$ and $\varepsilon=1/\sqrt{t}$. Here we show that, following a similar but simpler argument, the same limit results can be obtained for the linearized process.
3.2.1 Limit distribution of the linearized process
In this section, we show that the linearized process behaves similarly in the limit $t\to\infty$, and, in particular, that the invariance principle and central limit theorem (3.57) and (3.54), respectively, are also valid for the linearized process, and with the same covariance matrix for the limiting Brownian motion. We make the appropriate assumptions on the drift and diffusion coefficients that ensure that the mean field process is geometrically ergodic and we consider the linearized process $Y_{t}$ around the unique invariant measure $\phi_{\infty}$
$\mathrm{d}Y_{t}=-\nabla V(Y_{t})\,\mathrm{d}t-(\nabla W*\phi_{\infty})(Y_{t})\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},$ (3.58)
and let $\widetilde{Y}_{t}^{\varepsilon}\coloneqq\varepsilon Y_{t/\varepsilon^{2}}$. In the next two theorems we first prove the CLT for $Y_{t}$, whose argument is inspired by the proof of [22, Lemma 2.3], and then compute the limit of $\widetilde{Y}_{t}^{\varepsilon}$ following the methodology of [47]. These results, which are analogous to equations (3.57) and (3.54), respectively, show that the limit distributions are the same as those for the nonlinear process.
Theorem 3.9.
Let $Y_{t}$ be the solution of the SDE (3.58). Then, it holds
$\lim_{t\to\infty}\frac{1}{\sqrt{t}}Y_{t}=\mathcal{N}(0,2\beta^{-1}\mathcal{D}),\qquad\text{in law},$ (3.59)
where $\mathcal{D}$ is defined in equation (3.55).
Proof.
Let $\Phi$ be given by equation (3.56). Applying Itô’s lemma, we have
$\Phi(Y_{t})$ $=\Phi(Y_{0})+\sqrt{2\beta^{-1}}\int_{0}^{t}\nabla\Phi(Y_{s})\,\mathrm{d}B_{s}$ (3.60)
$\quad-\int_{0}^{t}\nabla\Phi(Y_{s})\left(\nabla V(Y_{s})+(\nabla W*\phi_{\infty})(Y_{s})\right)\,\mathrm{d}s+\beta^{-1}\int_{0}^{t}\Delta\Phi(Y_{s})\,\mathrm{d}s,$
and by equation (3.56) we obtain
$-\int_{0}^{t}\left(\nabla V(Y_{s})+(\nabla W*\phi_{\infty})(Y_{s})\right)\,\mathrm{d}s=\Phi(Y_{0})-\Phi(Y_{t})+\sqrt{2\beta^{-1}}\int_{0}^{t}\nabla\Phi(Y_{s})\,\mathrm{d}B_{s}.$ (3.61)
Therefore, we get
$Y_{t}$ $=Y_{0}-\int_{0}^{t}\left(\nabla V(Y_{s})+(\nabla W*\phi_{\infty})(Y_{s})\right)\,\mathrm{d}s+\sqrt{2\beta^{-1}}\int_{0}^{t}\mathrm{d}B_{s}$ (3.62)
$=Y_{0}+\Phi(Y_{0})-\Phi(Y_{t})+\sqrt{2\beta^{-1}}\int_{0}^{t}(I+\nabla\Phi(Y_{s}))\,\mathrm{d}B_{s},$
which implies
$\frac{1}{\sqrt{t}}Y_{t}=\frac{Y_{0}+\Phi(Y_{0})-\Phi(Y_{t})}{\sqrt{t}}+\sqrt{2\beta^{-1}}\frac{1}{\sqrt{t}}\int_{0}^{t}(I+\nabla\Phi(Y_{s}))\,\mathrm{d}B_{s}.$ (3.63)
Due to the boundedness of $\Phi$, we deduce that the first term in the right-hand side vanishes almost surely, and therefore in probability. Moreover, by the periodicity of $\Phi$ and the ergodic theorem we have
$\lim_{t\to\infty}\frac{1}{t}\int_{0}^{t}(I+\nabla\Phi(Y_{s}))(I+\nabla\Phi(Y_{s}))^{\top}\,\mathrm{d}s=\operatorname{\mathbb{E}}^{\phi_{\infty}}[(I+\nabla\Phi(Y))(I+\nabla\Phi(Y))^{\top}]=\mathcal{D}.$ (3.64)
Finally, by the martingale central limit theorem we deduce that
$\lim_{t\to\infty}\frac{1}{\sqrt{t}}\int_{0}^{t}(I+\nabla\Phi(Y_{s}))\,\mathrm{d}s=\mathcal{N}(0,\mathcal{D}),\qquad\text{in law},$ (3.65)
and the desired result follows by Slutsky’s theorem. ∎
Theorem 3.10.
Let $Y_{t}$ be the solution of the SDE (3.58), and let $\widetilde{Y}^{\varepsilon}_{t}=\varepsilon Y^{\varepsilon}_{t/\varepsilon^{2}}$. Then, it holds
$\lim_{\varepsilon\to 0}\widetilde{Y}^{\varepsilon}_{t}=\sqrt{2\beta^{-1}\mathcal{D}}B_{t},\qquad\text{in law},$ (3.66)
where $\mathcal{D}$ is defined in equation (3.55) and $B_{t}$ is a standard $d$-dimensional Brownian motion.
Proof.
Reasoning as in the proof of Theorem 3.9, we obtain
$\widetilde{Y}^{\varepsilon}_{t}=\varepsilon Y_{t/\varepsilon^{2}}=\varepsilon\left(Y_{0}+\Phi(Y_{0})-\Phi(Y_{t/\varepsilon^{2}})\right)+\varepsilon\sqrt{2\beta^{-1}}\int_{0}^{t/\varepsilon^{2}}(I+\nabla\Phi(Y_{s}))\,\mathrm{d}B_{s},$ (3.67)
where the first term in the right-hand side vanishes almost surely, and therefore in probability, as $\varepsilon\to 0$ due to the boundedness of $\Phi$. Hence, the limiting properties of the rescaled process $\widetilde{Y}^{\varepsilon}_{t}$ are determined studying the second part. Introducing the martingale
$\mathcal{M}_{\tau}=\sqrt{2\beta^{-1}}\int_{0}^{\tau}(I+\nabla\Phi(Y_{s}))\,\mathrm{d}B_{s},$ (3.68)
we can write
$\widetilde{Y}^{\varepsilon}_{t}=\varepsilon\left(Y_{0}+\Phi(Y_{0})-\Phi(Y_{t/\varepsilon^{2}})\right)+\varepsilon\mathcal{M}_{t/\varepsilon^{2}}.$ (3.69)
By the ergodic theorem and the periodicity of $\Phi$, the quadratic variation $[\mathcal{M}]_{t/\varepsilon^{2}}$ of the martingale term satisfies
$\lim_{\varepsilon\to 0}[\varepsilon\mathcal{M}]_{t/\varepsilon^{2}}=2\beta^{-1}t\lim_{\varepsilon\to 0}\frac{\varepsilon^{2}}{t}\int_{0}^{t/\varepsilon^{2}}(I+\nabla\Phi(Y_{s}))(I+\nabla\Phi(Y_{s}))^{\top}\,\mathrm{d}s=2\beta^{-1}t\mathcal{D}.$ (3.70)
We now define the difference martingale
$\mathfrak{M}_{s,t}^{\varepsilon}=\varepsilon\mathcal{M}_{t/\varepsilon^{2}}-\varepsilon\mathcal{M}_{s/\varepsilon^{2}},$ (3.71)
whose quadratic variation, reasoning as in equation (3.70), satisfies
$\lim_{\varepsilon\to 0}[\mathfrak{M}^{\varepsilon}]_{s,t}=2\beta^{-1}\lim_{\varepsilon\to 0}\varepsilon^{2}\int_{s/\varepsilon^{2}}^{t/\varepsilon^{2}}(I+\nabla\Phi(Y_{\tau}))(I+\nabla\Phi(Y_{\tau}))^{\top}\,\mathrm{d}\tau=2\beta^{-1}(t-s)\mathcal{D}.$ (3.72)
Finally, the desired result is obtained following the proof of [47, Theorem 1.5]. ∎
3.2.2 Numerical example
Figure 2: Numerical verification of the CLTs in equation (3.57) and Theorem 3.9, considering sinusoidal confining and interaction potentials.
We numerically verify in a one-dimensional test case the CLTs for the nonlinear and linear processes $X_{t}$ and $Y_{t}$ given in equations (3.57) and Theorem 3.9, respectively. We consider the setting of [22, Lemma 1.26] below the phase transition, i.e., we set $W(x)=-\cos(2\pi x)$, $V(x)=-\xi\cos(2\pi x)$ with $\xi=1/2$, and $\beta=1$. In [22, Lemma 1.26], it is proved that the wrapping of the solution of the McKean SDE has a unique invariant measure with density
$\phi_{\infty}(x)=\frac{e^{A\cos(2\pi x)}}{\int_{\mathbb{T}}e^{A\cos(2\pi z)}\,\mathrm{d}z},$ (3.73)
where $A$ is the solution of the nonlinear equation
$A=\beta\left(\xi+\frac{I_{1}(A)}{I_{0}(A)}\right),$ (3.74)
with $I_{0}$ and $I_{1}$ being modified Bessel functions of the first kind. The calculation of the limit diffusion coefficient, based upon the solution of the Poison equation (3.56), is identical to what is done in the standard theory of periodic homogenization [49, Section 13.6.1]. Hence, we deduce that the limit coefficient $\mathcal{D}$ is given by
$\mathcal{D}=\frac{1}{\left(\int_{\mathbb{T}}e^{A\cos(2\pi z)}\,\mathrm{d}z\right)^{2}}=I_{0}(A)^{-2}.$ (3.75)
Moreover, equation (3.58) for the linearized process $Y_{t}$ can be rewritten as
$\,\mathrm{d}Y_{t}=-2\pi\beta^{-1}A\sin(2\pi Y_{t})\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t}.$ (3.76)
In Figure 2 we show the histograms of $X_{t}/\sqrt{t}$ and $Y_{t}/\sqrt{t}$ with $t=10^{3}$ for $500$ different realizations of the Brownian motion. We remark that, differently from equation (3.76) that can be simulated, the McKean SDE for $X_{t}$ must be approximated by the interacting particle system (1.1), for which we use $N=250$ particles. All the SDEs have initial distribution $\mu_{0}=\nu_{0}=\delta_{0}$, and they are discretized using the Euler–Maruyama method with a time step $\Delta t=10^{-2}$. We observe that the histograms match the limit Gaussian distribution $\mathcal{N}(0,2\beta^{-1}\mathcal{D})$ predicted by the theory.
4 Conclusion
In this work, we considered the long-time behaviour of the McKean SDE, obtained in the mean field limit of a system of weakly interacting diffusions. We showed that, under the appropriate assumptions on the confining and interaction potentials that ensure that the McKean SDE is geometrically ergodic, we can “linearize” the nonlinear process in a systematic way and derive an Itô diffusion process that behaves like the nonlinear process at long times. This linear (in the sense of McKean) SDE is obtained by replacing the law of the process with the invariant measure of the dynamics. We considered the problem both with $\mathbb{R}^{d}$ and $\mathbb{T}^{d}$ as state spaces, and studied the asymptotic behavior of the law of the linearized process. We showed its exponentially fast convergence in time to the law of the nonlinear process, even for different initial distributions, both in relative entropy and Wasserstein distance. The rate of convergence depends on the coefficients of the McKean SDE via the LSI constant of the corresponding Gibbs measure and on the LSI of the initial distribution. Our analysis strongly relies on LSIs for the law of the linearized process. Moreover, we also proved strong convergence of the paths in $L^{2}$ when the state space is $\mathbb{R}^{d}$. We remark that an additional contribution of our work is the study of the convergence to equilibrium for the Mckean SDE on the torus with nonzero confining potential. In fact, under technical assumptions on the coefficients, we showed exponentially fast convergence to the invariant measure both in relative entropy and $L^{1}$ and $L^{2}$ for the densities.
Furthermore, we applied the linearization procedure to the study of the MLE for estimating unknown parameters in the McKean SDE. In particular, we considered the likelihood function of the linearized process, and proposed a “linearized” MLE that remains asymptotically unbiased in the limit of infinite data. We finally considered the combined diffusive-mean field limit for interacting particle systems in periodic interaction and confining potentials, and we verified that the limit distributions of the linear and nonlinear processes are the same.
In the same spirit as we did in [51], where we implicitly used this approach, as future developments of this work, we plan to use the techniques developed in this paper to analyze nonparametric estimators for interacting particle systems. A crucial component of this program is the accurate approximation of the invariant measure, which is used to derive the linearized McKean SDE.
The proposed methodology can be applied to other problems. We are particularly interested in the case where there are multiple invariant measures. We believe that our methodology can be applied to study the problem considered in [44]. In addition, we would like to obtain similar results for the underdamped/kinetic Mckean SDE and for the corresponding McKean–Vlasov PDE, as well as for coupled Vlasov–Poisson–Fokker–Planck systems. Another application of the proposed linearization procedure is the rigorous analysis of optimal control methodologies for the mean field dynamics that were developed in [5]. All of these topics are currently under investigation.
Acknowledgements
[…]

### v2

@@
 In this article, we consider McKean stochastic differential equations, as well as their corresponding McKean–Vlasov partial differential equations, which admit a unique stationary state, and we study the linearized Itô diffusion process that is obtained by replacing the law of the process in the convolution term with the unique invariant measure.
-We show that the law of this linearized process converges to the law of the nonlinear process exponentially fast in time, both in relative entropy and in Wasserstein distance.
+We show that the law of the nonlinear McKean process converges to the law of this linearized process exponentially fast in time, both in relative entropy and in Wasserstein distance.
 We study the problem in both the whole space and the torus.
@@
 AMS subject classifications. 26D10, 35Q70, 35Q83, 60J60, 82C22.
-Key words. Diffusive limit, interacting particle systems, invariant measure, logarithmic Sobolev inequality, maximum likelihood estimation, McKean SDE, McKean–Vlasov PDE, relative entropy.
+Key words. Diffusive limit, interacting particle systems, invariant measure, logarithmic Sobolev inequality, maximum likelihood estimation, McKean–Vlasov PDE.
 1 Introduction
@@
 Furthermore, the covariance matrix of the limiting Brownian motion can be calculated using the standard homogenization (Kipnis–Varadhan) formula [49, Section 13.6.1], for a linear elliptic cell problem, obtained by calculating the convolution in the generator of the McKean SDE with respect to the Gibbs measure of the dynamics in the torus.
-In the present work, we show that the the diffusive limit of the interacting particle system and the formula for the diffusion coefficient can be obtained from the linearization procedure.
+In the present work, we show that the diffusive limit of the interacting particle system and the formula for the diffusion coefficient can be obtained from the linearization procedure.
 In this paper we consider the dynamics of a system of $N$ weakly interacting reversible Langevin diffusions, described by the following system of Itô SDEs
@@
 $\mathcal{H}(\mu_{t}|\nu_{t})+\frac{1}{2\beta\lambda}\int_{0}^{t}\mathcal{H}(\mu_{s}|\nu_{s})\,\mathrm{d}s\leq\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta}{2}\int_{0}^{t}\int_{\mathbb{D}}\left\|(\nabla W*(\mu_{s}-\mu_{\infty}))(x)\right\|^{2}\mu_{s}(x)\,\mathrm{d}x\,\mathrm{d}s,$ (1.7)
-for some appropriate constant $\lambda>0$. The details can be found in Section 2. The fact that the laws $\mu_{t}$ and $\nu_{t}$ are exponentially close follows from this estimate.
+for some appropriate constant $\lambda>0$. The details can be found in Section˜2. The fact that the laws $\mu_{t}$ and $\nu_{t}$ are exponentially close follows from this estimate.
 We consider the problem both in $\mathbb{R}^{d}$ and in the torus $\mathbb{T}^{d}$.
@@
 The main contributions of this work can be summarized as follows.
-  • We prove that, under assumptions on the drift and confining potential that ensure uniform propagation of chaos, the McKean SDE on $\mathbb{R}^{d}$ is exponentially close in time to the linearized process, both in relative entropy and in the $2$-Wasserstein distance.
+  • We prove that, under assumptions on the confining and interaction potentials that ensure uniform propagation of chaos, the McKean SDE on $\mathbb{R}^{d}$ is exponentially close in time to the linearized process, both in relative entropy and in the 222-Wasserstein distance.
   • We obtain a similar result for the McKean SDE on the torus in relative entropy.
@@
 The remainder of the paper is organized as follows.
-In Section 2 we show the convergence analysis for the law of the linearized McKean SDE, which we perform both in the whole space in Section 2.2 and on the torus in Section 2.3.
-Then, in Section 3 we present two applications of the linearized process: maximum likelihood estimation in Section 3.1 and computation of the asymptotic variance in a central limit theorem in Section 3.2.
-Finally, in Section 4 we draw our conclusions and outline future developments of the linearization procedure.
+In Section˜2 we show the convergence analysis for the law of the linearized McKean SDE, which we perform both in the whole space in Section˜2.2 and on the torus in Section˜2.3.
+Then, in Section˜3 we present two applications of the linearized process: maximum likelihood estimation in Section˜3.1 and computation of the asymptotic variance in a central limit theorem in Section˜3.2.
+Finally, in Section˜4 we draw our conclusions and outline future developments of the linearization procedure.
 2 The linearized diffusion process
@@
 The situation is more complicated when we do not have uniqueness of the invariant measure.
-However, Assumption 2.1 is not a strong limitation of the methodology introduced in this work.
+However, Assumption˜2.1 is not a strong limitation of the methodology introduced in this work.
 In fact, in [44] it is shown that when multiple invariant measures can exist, it is still possible to prove exponentially fast convergence to equilibrium locally, i.e., as long as the initial distribution is sufficiently close to the stationary state and we are away from the critical temperature, where we only get algebraic convergence.
@@
 Lemma 2.4.
-Under Assumption 2.3, there exists $M>0$ such that for all $t\geq 0$
-$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left\lvert\nabla W(x-z)\right\rvert^{2}(f_{s}(z)+f_{\infty}(z))f_{s}(x)\,\mathrm{d}z\,\mathrm{d}x\leq M.$ (2.9)
+Under Assumption˜2.3, there exists $M>0$ such that for all $t\geq 0$
+$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left\lvert\nabla W(x-z)\right\rvert^{2}(f_{t}(z)+f_{\infty}(z))f_{t}(x)\,\mathrm{d}z\,\mathrm{d}x\leq M.$ (2.9)
 Proof.
@@
 Lemma 2.5.
-Let $W$ satisfy Assumption 2.3(iii), then, for all probability density functions $f$ of measures in $\mathcal{P}(\mathbb{R}^{d})$, the function $W*f$ is uniformly convex with constant $\gamma\geq 0$, i.e., it holds for all $x,y\in\mathbb{R}^{d}$
+Let $W$ satisfy Assumption˜2.3(iii), then, for all probability density functions $f$ of measures in $\mathcal{P}(\mathbb{R}^{d})$, the function $W*f$ is uniformly convex with constant $\gamma\geq 0$, i.e., it holds for all $x,y\in\mathbb{R}^{d}$
 $((\nabla W*f)(x)-(\nabla W*f)(y))^{\top}(x-y)\geq\gamma\left\lvert x-y\right\rvert^{2}.$ (2.10)
@@
 $\qquad=\int_{\mathbb{R}^{d}}(\nabla W(x-z)-\nabla W(y-z))^{\top}((x-z)-(y-z))f(z)\,\mathrm{d}z,$
-which due to Assumption 2.3(iii) implies
+which due to Assumption˜2.3(iii) implies
 $((\nabla W*f)(x)-(\nabla W*f)(y))^{\top}(x-y)\geq\int_{\mathbb{R}^{d}}\gamma\left\lvert x-y\right\rvert^{2}f(z)\,\mathrm{d}z=\gamma\left\lvert x-y\right\rvert^{2},$ (2.12)
@@
 2.2.1 Convergence of the linearized process
-In the classical work [41] as well in the interesting recent paper [43], it is shown that, provided that the initial condition satisfies an LSI, and under our convexity assumptions on the confining and interaction potentials given in Assumption 2.3, then so does the solution of the McKean–Vlasov PDE (1.3), with an LSI constant that can be explicitly computed.
+In the classical work [41] as well in the interesting recent paper [43], it is shown that, provided that the initial condition satisfies an LSI, and under our convexity assumptions on the confining and interaction potentials given in Assumption˜2.3, then so does the solution of the McKean–Vlasov PDE (1.3), with an LSI constant that can be explicitly computed.
 In the following lemma we show that the same conclusion holds for the linearized McKean–Vlasov PDE, with the same LSI constant.
 Lemma 2.6.
-Let Assumption 2.3 be satisfied and assume that $\mu_{0}$ and $\nu_{0}$ satisfy an LSI with constant $\lambda_{0}>0$. Then $\mu_{t}$ and $\nu_{t}$ satisfy an LSI with constant
+Let Assumption˜2.3 be satisfied and assume that $\mu_{0}$ and $\nu_{0}$ satisfy an LSI with constant $\lambda_{0}>0$. Then $\mu_{t}$ and $\nu_{t}$ satisfy an LSI with constant
 $\lambda_{t}=\lambda_{0}e^{-2(\alpha+\gamma)t}+\frac{1}{2\beta(\alpha+\gamma)}\left(1-e^{-2(\alpha+\gamma)t}\right).$ (2.13)
@@
 Proof.
-The desired result follows directly from [43, Proposition 1.1] and Lemma 2.5, and the uniform in time constant is obtained by noticing that $\lambda_{t}$ is a convex combination of $\lambda_{0}$ and $1/(2\beta(\alpha+\gamma))$.
+The desired result follows directly from [43, Proposition 1.1] and Lemma˜2.5, and the uniform in time constant is obtained by noticing that $\lambda_{t}$ is a convex combination of $\lambda_{0}$ and $1/(2\beta(\alpha+\gamma))$.
 ∎
@@
 Theorem 2.7.
-Let the assumptions of Lemma 2.6 be satisfied. Then, it holds
+Let the assumptions of Lemma˜2.6 be satisfied. Then, it holds
 $\mathcal{H}(\mu_{t}|\nu_{t})\leq\rho_{\Lambda}(t)\coloneqq\begin{cases}\left(\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta^{2}MK\Lambda}{4-\alpha\beta\Lambda}\right)e^{-\frac{1}{2}\alpha t},&\text{if }\Lambda<\frac{4}{\alpha\beta},\\ \left(\mathcal{H}(\mu_{0}|\nu_{0})+\frac{1}{2}\beta MKt\right)e^{-\frac{2}{\beta\Lambda}t},&\text{if }\Lambda=\frac{4}{\alpha\beta},\\ \left(\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta^{2}MK\Lambda}{\alpha\beta\Lambda-4}\right)e^{-\frac{2}{\beta\Lambda}t},&\text{if }\Lambda>\frac{4}{\alpha\beta}.\end{cases}$ (2.15)
@@
 $\mathcal{H}(\mu_{t}|\nu_{t})+\frac{1}{2\beta}\int_{0}^{t}\mathcal{I}(\mu_{s}|\nu_{s})\,\mathrm{d}s\leq\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta}{2}\int_{0}^{t}\int_{\mathbb{R}^{d}}\left\lvert(\nabla W*(f_{s}-f_{\infty}))(x)\right\rvert^{2}f_{s}(x)\,\mathrm{d}x\,\mathrm{d}s,$ (2.16)
-which due to Lemma 2.6 implies
+which due to Lemma˜2.6 implies
 $\mathcal{H}(\mu_{t}|\nu_{t})+\frac{2}{\beta\Lambda}\int_{0}^{t}\mathcal{H}(\mu_{s}|\nu_{s})\,\mathrm{d}s\leq\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta}{2}\int_{0}^{t}\int_{\mathbb{R}^{d}}\left\lvert(\nabla W*(f_{s}-f_{\infty}))(x)\right\rvert^{2}f_{s}(x)\,\mathrm{d}x\,\mathrm{d}s.$ (2.17)
@@
 $\mathcal{H}(\mu_{t}|\nu_{t})\leq\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta MK}{2}\int_{0}^{t}e^{-\frac{1}{2}\alpha s}\,\mathrm{d}s-\frac{2}{\beta\Lambda}\int_{0}^{t}\mathcal{H}(\mu_{s}|\nu_{s})\,\mathrm{d}s,$ (2.20)
-which, by [39, Lemma A.1], implies
+which, by the Grönwall’s inequality in [39, Lemma A.1], implies
 $\mathcal{H}(\mu_{t}|\nu_{t})$ $\leq\mathcal{H}(\mu_{0}|\nu_{0})e^{-\frac{2}{\beta\Lambda}t}+\frac{\beta MK}{2}\int_{0}^{t}e^{-\frac{2}{\beta\Lambda}(t-s)}e^{-\frac{1}{2}\alpha s}\,\mathrm{d}s,$ (2.21)
@@
 Corollary 2.8.
-Let the assumptions of Theorem 2.7 be satisfied. Then, it holds
+Let the assumptions of Theorem˜2.7 be satisfied. Then, it holds
 $\mathcal{W}_{2}(\mu_{t},\nu_{t})\leq\sqrt{\Lambda\rho_{\Lambda}(t)},$ (2.24)
-where $\mathcal{W}_{2}$ denotes the 2-Wasserstein distance and $\rho_{\Lambda}(t)$ is defined in Theorem 2.7.
+where $\mathcal{W}_{2}$ denotes the 2-Wasserstein distance and $\rho_{\Lambda}(t)$ is defined in Theorem˜2.7.
 Proof.
-The desired result follows from Theorem 2.7 using [48, Theorem 1]. ∎
+The desired result follows from Theorem˜2.7 using [48, Theorem 1]. ∎
 Finally, we also have strong convergence of the paths $X_{t}$ and $Y_{t}$ in $L^{2}$.
 Theorem 2.9.
-Let Assumption 2.3 be satisfied. Then, it holds
+Let Assumption˜2.3 be satisfied. Then, it holds
 $\operatorname{\mathbb{E}}[\left\|X_{t}-Y_{t}\right\|^{2}]\leq\left(\operatorname{\mathbb{E}}[\left\|X_{0}-Y_{0}\right\|^{2}]+\frac{8MK}{(3\alpha+4\gamma)^{2}}\right)e^{-\frac{1}{2}\alpha t}.$ (2.25)
@@
 $\eqqcolon I_{2,1}+I_{2,2},$
-where Lemma 2.5 gives
+where Lemma˜2.5 gives
 $I_{2,2}\leq-\gamma\left\lvert X_{t}-Y_{t}\right\rvert^{2}.$ (2.30)
@@
 Remark 2.10.
-If the LSI constant $\lambda_{0}$ of the initial distributions $\mu_{0}$ and $\nu_{0}$ is sufficiently small, namely $\lambda_{0}<4/(\alpha\beta)$, then the rate of convergence in Theorems 2.7 and 2.8 is determined only by the convexity parameter $\alpha$ of the confining potential.
-This is in agreement with the rate of convergence in $L^{2}$ predicted by Theorem 2.9.
+If the LSI constant $\lambda_{0}$ of the initial distributions $\mu_{0}$ and $\nu_{0}$ is sufficiently small, namely $\lambda_{0}<4/(\alpha\beta)$, then the rate of convergence in Theorems˜2.7 and 2.8 is determined only by the convexity parameter $\alpha$ of the confining potential.
+This is in agreement with the rate of convergence in $L^{2}$ predicted by Theorem˜2.9.
 2.3 Analysis on $\mathbb{T}^{d}$
-In this section, the domain $\mathbb{D}$ is the multidimensional torus $\mathbb{T}^{d}$ of size length $1$, and we work in the framework below.
+In this section, the domain $\mathbb{D}$ is the multidimensional torus $\mathbb{T}^{d}$ of size length 111, and we work in the framework below.
 Assumption 2.11.
@@
     W is either $H$-stable, meaning that it has nonnegative Fourier modes [55], or is such that $\left\|W\right\|_{L^{\infty}(\mathbb{T}^{d})}<\beta^{-1}$.
-Notice that Assumption 2.11(iii) guarantees that Assumption 2.1 is satisfied. In fact, all invariant probability density functions $f_{\infty}$ are also minimizers of the free energy
+Notice that Assumption˜2.11(iii) guarantees that Assumption˜2.1 is satisfied. In fact, all invariant probability density functions $f_{\infty}$ are also minimizers of the free energy
 $\mathcal{E}(f)=\int_{\mathbb{T}^{d}}V(x)f(x)\,\mathrm{d}x+\frac{1}{2}\int_{\mathbb{T}^{d}}\int_{\mathbb{T}^{d}}W(x-y)f(y)f(x)\,\mathrm{d}y\,\mathrm{d}x+\beta^{-1}\int_{\mathbb{T}^{d}}f(x)\log(f(x))\,\mathrm{d}x,$ (2.35)
-which is strictly convex under Assumption 2.11(iii).
+which is strictly convex under Assumption˜2.11(iii).
 This is shown in [14, Proposition 2.8] without the confining potential, which, however, does not affect the convexity of $\mathcal{E}$ since its corresponding term is linear; see also [3, Section V.B.].
@@
 Lemma 2.12.
-Under Assumption 2.11, the invariant measure $\mathrm{d}\mu_{\infty}(x)=f_{\infty}(x)\,\mathrm{d}x$ satisfies an LSI with constant $\Gamma/(2\pi^{2})$, where
+Under Assumption˜2.11, the invariant measure $\mathrm{d}\mu_{\infty}(x)=f_{\infty}(x)\,\mathrm{d}x$ satisfies an LSI with constant $\Gamma/(2\pi^{2})$, where
 $\Gamma=e^{2\beta\left(\left\|V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|W\right\|_{L^{\infty}(\mathbb{T}^{d})}\right)}.$ (2.36)
 Proof.
-We follow [33, Lemma 2.4]. From the integral equation (1.4) we deduce that for all $x\in\mathbb{T}^{d}$
+We follow [33, Lemma 2.4]. From the integral equation (1.4) and the Young’s convolution inequality we deduce that for all $x\in\mathbb{T}^{d}$
+$f_{\infty}(x)=\frac{e^{-\beta(V(x)+(W*f_{\infty})(x))}}{\int_{\mathbb{T}^{d}}e^{-\beta(V(z)+(W*f_{\infty})(z))}\,\mathrm{d}z}\leq\frac{e^{\beta\left(\left\|V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|W\right\|_{L^{\infty}(\mathbb{T}^{d})}\left\|f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}\right)}}{\int_{\mathbb{T}^{d}}e^{-\beta\left(\left\|V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|W\right\|_{L^{\infty}(\mathbb{T}^{d})}\left\|f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}\right)}\,\mathrm{d}z}.$ (2.37)
+Proceeding analogously for the lower bound, and noting that $\left\|f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}=1$ since $f_{\infty}$ is a probability density function and that $\mathbb{T}^{d}$ has Lebesgue measure 1, we obtain
 $\frac{1}{\Gamma}=e^{-2\beta\left(\left\|V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|W\right\|_{L^{\infty}(\mathbb{T}^{d})}\right)}\leq f_{\infty}(x)\leq e^{2\beta\left(\left\|V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|W\right\|_{L^{\infty}(\mathbb{T}^{d})}\right)}=\Gamma.$ (2.38)
 Then, the desired result follows from the Holley–-Stroock perturbation lemma (see, e.g., [33, Lemma 2.3]) and the fact that the uniform distribution on the torus satisfies an LSI (see, e.g., [33, Lemma 2.4]).
-We remark that the discrepancy with [33, Lemma 2.4] is due to the factor $4$ in the LSI (2.6).
+We remark that the discrepancy with [33, Lemma 2.4] is due to the factor 444 in the LSI (2.6).
 ∎
@@
 This problem has already been studied when there is no confining potential in, e.g., [17, 14].
-In the absence of confining potential, the analysis is simplified because the unique, under Assumption 2.11(iii), invariant measure is the Lebesgue measure.
+In the absence of confining potential, the analysis is simplified because the unique, under Assumption˜2.11(iii), invariant measure is the Lebesgue measure.
 However, here we extend the results to the case of a nonzero confining potential.
 Given two probabilities $\mu,\nu\in\mathcal{P}(\mathbb{T}^{d})$ with densities $f,g$, we denote their total variation distance by $\mathrm{TV}(\mu,\nu)$.
-Moreover, we recall that it holds the following equivalence with the $L^{1}(\mathbb{T}^{d})$ distance between the densities
+Moreover, we recall the following equivalence with the $L^{1}(\mathbb{T}^{d})$ distance between the densities (see, e.g., [61, Lemma 2.1])
 $\mathrm{TV}(\mu,\nu)=\frac{1}{2}\left\|f-g\right\|_{L^{1}(\mathbb{T}^{d})}.$ (2.39)
@@
 Proposition 2.14.
-Let Assumption 2.11 be satisfied. Then, it holds
+Let Assumption˜2.11 be satisfied. Then, it holds
 $\left\|f_{t}-f_{\infty}\right\|_{L^{2}(\mathbb{T}^{d})}\leq\left\|f_{0}-f_{\infty}\right\|_{L^{2}(\mathbb{T}^{d})}e^{-\zeta t},$ (2.40)
@@
 which, due to Hölder’s and Young’s convolution inequalities, implies
-$\left\lvert I_{t}\right\rvert$ $\leq\left(\left\|\nabla V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}\right)\left\|f_{t}-f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}\left\|\nabla(f_{t}-f_{\infty})\right\|_{L^{1}(\mathbb{T}^{d})}$ (2.43)
+$\left\lvert I_{t}\right\rvert$ $\leq\left(\left\|\nabla V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}\right)\left\|f_{t}-f_{\infty}\right\|_{L^{2}(\mathbb{T}^{d})}\left\|\nabla(f_{t}-f_{\infty})\right\|_{L^{2}(\mathbb{T}^{d})}$ (2.44)
 $\quad+\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}\left\|f_{\infty}\right\|_{L^{\infty}(\mathbb{T}^{d})}\left\|f_{t}-f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}\left\|\nabla(f_{t}-f_{\infty})\right\|_{L^{1}(\mathbb{T}^{d})}.$
@@
 Proposition 2.15.
-Let Assumption 2.11 be satisfied. Then, it holds
+Let Assumption˜2.11 be satisfied. Then, it holds
 $\mathcal{H}(f_{t}|f_{\infty})\leq\mathcal{H}(f_{0}|f_{\infty})e^{-\eta t},$ (2.49)
@@
 $\eqqcolon-\beta^{-1}\mathcal{I}(f_{t}|f_{\infty})+I^{(1)}_{t}+I^{(2)}_{t}.$
-We now need to bound the terms in the right-hand side. First, by Hölder’s and Young’s convolution inequalities, we have
+We now need to bound the terms in the right-hand side. First, by Hölder’s and Young’s convolution inequalities and since $\nabla\log f_{\infty}=-\beta(\nabla V+\nabla W*f_{\infty})$, we have
 $\left\lvert I^{(1)}_{t}\right\rvert$ $\leq\left\|\Delta W\right\|_{L^{\infty}(\mathbb{T}^{d})}\left\|f_{t}-f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}^{2},$ (2.57)
@@
 $\left\|f_{t}-f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}^{2}\leq 8\mathcal{H}(f_{t}|f_{\infty}).$ (2.58)
-Finally, Lemma 2.12 yields
+Finally, Lemma˜2.12 yields
 $\mathcal{I}(f_{t}|f_{\infty})\geq\frac{8\pi^{2}}{\Gamma}\mathcal{H}(f_{t}|f_{\infty}),$ (2.59)
@@
 Remark 2.16.
-In Propositions 2.14 and 2.15, the exponents $\zeta$ and $\eta$ are always strictly positive in the high temperature regime, i.e., when $\beta$ is sufficiently large.
-Moreover, from the proof of [13, Proposition 3.1] it follows that we can replace $\left\|\Delta W\right\|_{L^{\infty}(\mathbb{T}^{d})}$ in (2.49) with $\left\|\Delta W_{u}\right\|_{L^{\infty}(\mathbb{T}^{d})}$, where $W_{u}$ denotes the non-$H$-stable part of $W$ and is defined in [13, Definition 2.1].
-However, the same argument cannot be applied to $\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}$.
-The corollary below is a straightforward consequence of Propositions 2.14 and 2.15, and quantifies the rate of convergence in $L^{1}(\mathbb{T}^{d})$.
+[…]
+The corollary below is a straightforward consequence of Propositions˜2.14 and 2.15, and quantifies the rate of convergence in $L^{1}(\mathbb{T}^{d})$.
 Corollary 2.17.
-Let Assumption 2.11 be satisfied. Then, it holds for $i=1,2$
+Let Assumption˜2.11 be satisfied. Then, it holds for $i=1,2$
 $\left\|f_{t}-f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}\leq c_{i}e^{-a_{i}t},$ (2.61)
@@
 Proof.
-The desired result follows applying Propositions 2.14 and 2.15, the CKP inequality (2.57), and bounding the $L^{1}(\mathbb{T}^{d})$ norm with the $L^{2}(\mathbb{T}^{d})$ one. ∎
+The desired result follows applying Propositions˜2.14 and 2.15, the CKP inequality (2.58), and bounding the $L^{1}(\mathbb{T}^{d})$ norm with the $L^{2}(\mathbb{T}^{d})$ one. ∎
 2.3.2 Convergence of the linearized process
@@
 Lemma 2.18.
-Let Assumption 2.11 be satisfied, and assume that there exists a constant $\kappa>1$ such that for all $x\in\mathbb{T}^{d}$
+Let Assumption˜2.11 be satisfied, and assume that there exists a constant $\kappa>1$ such that for all $x\in\mathbb{T}^{d}$
 $\frac{1}{\kappa}\leq f_{0}(x)\leq\kappa\qquad\text{and}\qquad\frac{1}{\kappa}\leq g_{0}(x)\leq\kappa,$ (2.63)
@@
 where $\Gamma$ is defined in (2.36).
-Moreover, if $a_{i}>0$ and $C_{i}/a_{i}<\mathbb{W}(1)$ for either $i=1$ or $i=2$, with $a_{i}$ and $C_{i}$ given in Corollary 2.17 and equation (2.77) below, and where $\mathbb{W}$ denotes the Lambert function, then $\mu_{t}$ satisfies a time-uniform LSI with constant
+Moreover, if $a_{i}>0$ and $C_{i}/a_{i}<\mathbb{W}(1)$ for either $i=1$ or $i=2$, with $a_{i}$ and $C_{i}$ given in Corollary˜2.17 and equation (2.78) below, and where $\mathbb{W}$ denotes the Lambert function, then $\mu_{t}$ satisfies a time-uniform LSI with constant
 $\widetilde{\Xi}_{i}=\frac{\kappa\Gamma^{2}}{2\pi^{2}(1-(C_{i}/a_{i})e^{C_{i}/a_{i}})}.$ (2.65)
@@
 $\operatorname{\mathbb{E}}[\psi_{T-t}(Z_{t}^{\mu})]$ $=\psi_{T}(x)-\int_{0}^{t}\operatorname{\mathbb{E}}\left[\Delta W*((\psi_{T-s}-1)f_{\infty})(Z_{s}^{\mu})\psi_{T-s}(Z_{s}^{\mu})\,\mathrm{d}s\right]$ (2.75)
-$+\beta\int_{0}^{t}\operatorname{\mathbb{E}}\left[(\nabla V(Z_{s}^{\mu})+\nabla W*f_{\infty}(Z_{s}^{\mu}))\cdot\nabla W*((\psi_{T-s}-1)f_{\infty})(Z_{s}^{\mu})\psi_{T-s}(Z_{s}^{\mu})\,\mathrm{d}s\right].$
-Using Corollary 2.17 we have
+$\hskip-14.22636pt+\beta\int_{0}^{t}\operatorname{\mathbb{E}}\left[(\nabla V(Z_{s}^{\mu})+\nabla W*f_{\infty}(Z_{s}^{\mu}))\cdot\nabla W*((\psi_{T-s}-1)f_{\infty})(Z_{s}^{\mu})\psi_{T-s}(Z_{s}^{\mu})\,\mathrm{d}s\right].$
+Using Corollary˜2.17 we have
 $\left\|(\psi_{T-s}-1)f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}\leq c_{i}e^{-a_{i}(T-s)},$ (2.76)
@@
 $\kappa\Gamma e^{C_{i}/a_{i}}\leq\frac{\kappa\Gamma}{1-(C_{i}/a_{i})e^{C_{i}/a_{i}}}.$ (2.84)
-Therefore, the desired results follow from the bounds (2.71) and (2.82) by applying the Holley–Stroock perturbation lemma (see, e.g., [33, Lemma 2.3]) and Lemma 2.12. ∎
-Using the LSI, we can finally prove that the measures $\mu_{t}$ and $\nu_{t}$ are exponentially close in time, in relative entropy. The argument is similar to the proof of Theorem 2.7.
+Therefore, the desired results follow from the bounds (2.72) and (2.83) by applying the Holley–Stroock perturbation lemma (see, e.g., [33, Lemma 2.3]) and Lemma˜2.12. ∎
+Using the LSI, we can finally prove that the measures $\mu_{t}$ and $\nu_{t}$ are exponentially close in time, in relative entropy. The argument is similar to the proof of Theorem˜2.7.
 Theorem 2.19.
-Let the assumptions of Lemma 2.18 be satisfied. Then, it holds for $i=1,2$
+Let the assumptions of Lemma˜2.18 be satisfied. Then, it holds for $i=1,2$
 $\mathcal{H}(\mu_{t}|\nu_{t})\leq\sigma_{\Xi}(t)\coloneqq\begin{cases}\left(\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta^{2}c_{i}\Xi\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}^{2}}{2(2-\beta a_{i}\Xi)}\right)e^{-a_{i}t},&\text{if }\Xi<\frac{2}{a_{i}\beta},\\ \left(\mathcal{H}(\mu_{0}|\nu_{0})+\frac{1}{2}\beta c_{i}\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}^{2}t\right)e^{-\frac{2}{\beta\Xi}t},&\text{if }\Xi=\frac{2}{a_{i}\beta},\\ \left(\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta^{2}c_{i}\Xi\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}^{2}}{2(\beta a_{i}\Xi-2)}\right)e^{-\frac{2}{\beta\Xi}t},&\text{if }\Xi>\frac{2}{a_{i}\beta},\end{cases}$ (2.85)
-where $a_{i}$ and $c_{i}$ are defined in Corollary 2.17.
+where $a_{i}$ and $c_{i}$ are defined in Corollary˜2.17.
 Proof.
@@
 $\mathcal{H}(\mu_{t}|\nu_{t})+\frac{1}{2\beta}\int_{0}^{t}\mathcal{I}(\mu_{s}|\nu_{s})\,\mathrm{d}s\leq\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta}{2}\int_{0}^{t}\int_{\mathbb{T}^{d}}\left\lvert(\nabla W*(f_{s}-f_{\infty}))(x)\right\rvert^{2}f_{s}(x)\,\mathrm{d}x\,\mathrm{d}s,$ (2.86)
-which, due to Lemma 2.18, implies
+which, due to Lemma˜2.18, implies
 $\mathcal{H}(\mu_{t}|\nu_{t})+\frac{2}{\beta\Xi}\int_{0}^{t}\mathcal{H}(\mu_{s}|\nu_{s})\,\mathrm{d}s\leq\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta}{2}\int_{0}^{t}\int_{\mathbb{T}^{d}}\left\lvert(\nabla W*(f_{s}-f_{\infty}))(x)\right\rvert^{2}f_{s}(x)\,\mathrm{d}x\,\mathrm{d}s.$ (2.87)
-Then, Corollary 2.17 gives
+Then, Corollary˜2.17 gives
 $\left\lvert(\nabla W*(f_{s}-f_{\infty}))(x)\right\rvert^{2}\leq\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}^{2}\left\|f_{s}-f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}^{2}\leq c_{i}\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}^{2}e^{-a_{i}s},$ (2.88)
@@
 $\mathcal{H}(\mu_{t}|\nu_{t})\leq\mathcal{H}(\mu_{0}|\nu_{0})+\frac{\beta c_{i}\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}^{2}}{2}\int_{0}^{t}e^{-a_{i}s}\,\mathrm{d}s-\frac{2}{\beta\Xi}\int_{0}^{t}\mathcal{H}(\mu_{s}|\nu_{s})\,\mathrm{d}s.$ (2.89)
-Applying [39, Lemma A.1], we have
+Applying the Grönwall’s inequality in [39, Lemma A.1], we have
 $\mathcal{H}(\mu_{t}|\nu_{t})$ $\leq\mathcal{H}(\mu_{0}|\nu_{0})e^{-\frac{2}{\beta\Xi}t}+\frac{\beta c_{i}\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}^{2}}{2}\int_{0}^{t}e^{-\frac{2}{\beta\Xi}(t-s)}e^{-a_{i}s}\,\mathrm{d}s,$ (2.90)
@@
 Corollary 2.20.
-Let the assumptions of Theorem 2.19 be satisfied. Then, it holds
+Let the assumptions of Theorem˜2.19 be satisfied. Then, it holds
 $\mathcal{W}_{2}(\mu_{t},\nu_{t})\leq\sqrt{\Xi\sigma_{\Xi}(t)},$ (2.93)
-where $\mathcal{W}_{2}$ denotes the 2-Wasserstein distance and $\sigma_{\Xi}(t)$ is defined in Theorem 2.19.
+where $\mathcal{W}_{2}$ denotes the 2-Wasserstein distance and $\sigma_{\Xi}(t)$ is defined in Theorem˜2.19.
 Proof.
-The desired result follows from Theorem 2.19 using [48, Theorem 1]. ∎
+The desired result follows from Theorem˜2.19 using [48, Theorem 1]. ∎
 3 Applications
@@
 3.1 Maximum likelihood estimation
-We consider the case where the confining and interaction potentials $V$ and $W$ depend on unknown parameters.
-Our goal is to infer these parameters by observing a trajectory of the mean field SDE (1.2).
-Hence, consider the equation on the whole space $\mathbb{R}^{d}$ and in the time horizon $[0,T]$
+[…]
 $\mathrm{d}X_{t}=-\nabla V(X_{t};\theta)\,\mathrm{d}t-(\nabla W(\cdot;\theta)*f_{t}(\cdot;\theta))(X_{t})\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},$ (3.1)
@@
 $\widehat{\theta}_{T}=\operatorname{argmax}_{\theta\in\Theta}\mathfrak{L}_{T}(\theta)=\operatorname{argmax}_{\theta\in\Theta}\ell_{T}(\theta).$ (3.6)
-We now define the linearized version of the log-likelihood around the invariant measure $f_{\infty}$,
+We now define the linearized version of the log-likelihood around the invariant measure $f_{\infty}(\cdot;\theta)$,
 $\widetilde{\mathfrak{L}}_{T}(\theta)$ $=-\frac{\beta}{2}\int_{0}^{T}\left(\nabla V(X_{t};\theta)+(\nabla W(\cdot;\theta)*f_{\infty}(\cdot;\theta))(X_{t})\right)^{\top}\,\mathrm{d}X_{t}$ (3.7)
@@
     The function
-    $\theta\mapsto\int_{\mathbb{R}^{d}}\left\lvert\widetilde{b}_{\theta}(x)-\widetilde{b}_{\theta_{0}}(x)\right\rvert^{2}f_{\infty}(x;\theta)\,\mathrm{d}x$ (3.12)
+    $\theta\mapsto\int_{\mathbb{R}^{d}}\left\lvert\widetilde{b}_{\theta}(x)-\widetilde{b}_{\theta_{0}}(x)\right\rvert^{2}f_{\infty}(x;\theta_{0})\,\mathrm{d}x$ (3.12)
     admits only one zero at $\theta=\theta_{0}$.
 Remark 3.2.
-The condition in Assumption 3.1(iii) is necessary to guarantee the identifiability of the unknown parameter $\theta_{0}$, and it is related to the nondegeneracy of the Fisher information matrix in [24].
+The condition in Assumption˜3.1(iii), which appears also in [63], is necessary to guarantee the identifiability of the unknown parameter $\theta_{0}$, and it is related to the nondegeneracy of the Fisher information matrix in [24].
+We notice that this condition is satisfied, e.g., if $\widetilde{b}_{\theta}$ is an affine function of $\theta$.
 The main result of this section is the asymptotic unbiasedness of $\widetilde{\theta}_{T}$, when evaluated at a path of the McKean SDE.
 Theorem 3.3.
-Under Assumptions 2.3 and 3.1, it holds
+Under Assumptions˜2.3 and 3.1, it holds
 […]
@@
 In fact, in our approach the likelihood function is computed from the linear process, and the nonlinearity only appears in the observation model.
-The proof of Theorem 3.3 relies on the following preliminary results.
+The proof of Theorem˜3.3 relies on the following preliminary results.
 Lemma 3.5.
-Under Assumptions 2.3 and 3.1, it holds for all $\theta\in\Theta$
+Under Assumptions˜2.3 and 3.1, it holds for all $\theta\in\Theta$
 […]
@@
 Corollary 3.6.
-Under Assumptions 2.3 and 3.1, it holds
+Under Assumptions˜2.3 and 3.1, it holds
 […]
 Proof.
-The desired result follows from Lemma 3.5 by replacing $\theta$ with $\theta_{0}$. ∎
+The desired result follows from Lemma˜3.5 by replacing $\theta$ with $\theta_{0}$. ∎
 Lemma 3.7.
-Under Assumptions 2.3 and 3.1, it holds
+Under Assumptions˜2.3 and 3.1, it holds
 […]
@@
 $e(T,\theta)=\frac{1}{T}\int_{0}^{T}\left(\widetilde{b}_{\theta}(X_{t})-b_{\theta_{0}}(t,X_{t})\right)\,\mathrm{d}B_{t},$ (3.22)
-which is uniformly continuous in $\theta$ on $\Theta$, uniformly in $T\in[1,\infty)$, with probability $1$, by [11, Theorem 3.1].
-Then, by Lemma 3.5 and the strong law of large numbers for martingales we deduce that for all $\theta\in\Theta$
+which is uniformly continuous in $\theta$ on $\Theta$, uniformly in $T\in[1,\infty)$, with probability 111, by [11, Theorem 3.1].
+Then, by Lemma˜3.5 and the strong law of large numbers for martingales we deduce that for all $\theta\in\Theta$
 […]
@@
 Lemma 3.8.
-Under Assumptions 2.3 and 3.1, there exists a constant $L>0$, independent of $T$, such that for all $\theta_{1},\theta_{2}\in\Theta$
+Under Assumptions˜2.3 and 3.1, there exists a constant $L>0$, independent of $T$, such that for all $\theta_{1},\theta_{2}\in\Theta$
 $\left\lvert\frac{1}{T}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta_{1}}(X_{t})-b_{\theta_{0}}(t,X_{t})\right\rvert^{2}\,\mathrm{d}t-\frac{1}{T}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta_{2}}(X_{t})-b_{\theta_{0}}(t,X_{t})\right\rvert^{2}\,\mathrm{d}t\right\rvert\leq L\left\lvert\theta_{1}-\theta_{2}\right\rvert.$ (3.24)
@@
 $\hskip 42.67912pt\leq B\left(\frac{1}{T}\int_{0}^{T}\left\lvert\widetilde{b}_{\theta_{1}}(X_{t})-\widetilde{b}_{\theta_{2}}(X_{t})\right\rvert^{2}\right)^{1/2},$
-where the uniform bound in time of the last inequality is due to Lemma 3.5. By the definition of $\widetilde{b}_{\theta}$ in equation (3.4) we obtain
+where the uniform bound in time of the last inequality is due to Lemma˜3.5. By the definition of $\widetilde{b}_{\theta}$ in equation (3.4) we obtain
 $\left\lvert h(T,\theta_{1})-h(T,\theta_{2})\right\rvert^{2}$ $\leq\frac{B}{T}\int_{0}^{T}\left\lvert\nabla V(X_{t};\theta_{1})-\nabla V(X_{t};\theta_{2})\right\rvert^{2}\,\mathrm{d}t$ (3.27)
@@
 $+\frac{B}{T}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left\lvert\nabla W(X_{t}-y;\theta_{1})-\nabla W(X_{t}-y;\theta_{2})\right\rvert^{2}f_{\infty}(y;\theta_{2})\,\mathrm{d}y\,\mathrm{d}t.$
-Moreover, since $\nabla V$ and $\nabla W$ are Lipschitz in $\theta$ by Assumption 3.1(ii) and polynomially bounded in $x$, using the uniform boundedness of the moments and the ergodic theorem, and due to the differentiability of the invariant measure with respect to $\theta$ [9], we have that there exists a constant $L>0$, independent of $T$, such that
+Moreover, since $\nabla V$ and $\nabla W$ are Lipschitz in $\theta$ by Assumption˜3.1(ii) and polynomially bounded in $x$, using the uniform boundedness of the moments and the ergodic theorem, and due to the differentiability of the invariant measure with respect to $\theta$ [9], we have that there exists a constant $L>0$, independent of $T$, such that
 $\left\lvert h(T,\theta_{1})-h(T,\theta_{2})\right\rvert^{2}\leq L^{2}\left\lvert\theta_{1}-\theta_{2}\right\rvert^{2},$ (3.29)
 which yields the desired result. ∎
-We can finally prove Theorem 3.3.
-Proof of Theorem 3.3.
+We can finally prove Theorem˜3.3.
+Proof of Theorem˜3.3.
 By definition, the estimator $\widetilde{\theta}_{T}$ satisfies for all $\theta\in\Theta$
@@
 $\hskip 42.67912pt-\frac{\beta}{2T}\int_{0}^{T}\left(\widetilde{b}_{\theta_{0}}(X_{t})-b_{\theta_{0}}(t,X_{t})\right)^{\top}\,\mathrm{d}B_{t}\geq\frac{\beta}{4T}\int_{0}^{T}\left\lvert\widetilde{b}_{\widetilde{\theta}_{T}}(X_{t})-b_{\theta_{0}}(t,X_{t})\right\rvert^{2}\,\mathrm{d}t\geq 0.$
-Applying Corollaries 3.6 and 3.7 we obtain
+Applying Corollaries˜3.6 and 3.7 we obtain
 […]
-Let us now define the function $h=h(T,\theta)$ as in equation (3.25), which is Lipschitz in $\theta$, uniformly in $T$, by Lemma 3.8, and notice that
+Let us now define the function $h=h(T,\theta)$ as in equation (3.25), which is Lipschitz in $\theta$, uniformly in $T$, by Lemma˜3.8, and notice that
 […]
@@
 $\Omega=\left\{\theta\in\Theta\colon\lim_{T\to\infty}h(T,\theta)=0\right\},$ (3.35)
-which by definition (3.25) and Lemma 3.5 coincides with
-$\Omega=\left\{\theta\in\Theta\colon\int_{\mathbb{R}^{d}}\left\lvert\widetilde{b}_{\theta}(x)-\widetilde{b}_{\theta_{0}}(x)\right\rvert^{2}f_{\infty}(x)\,\mathrm{d}x=0\right\}.$ (3.36)
-Finally, by the identifiability condition in Assumption 3.1(iii) we get that $\Omega=\{\theta_{0}\}$, and the desired result follows. ∎
+which by definition (3.25) and Lemma˜3.5 coincides with
+$\Omega=\left\{\theta\in\Theta\colon\int_{\mathbb{R}^{d}}\left\lvert\widetilde{b}_{\theta}(x)-\widetilde{b}_{\theta_{0}}(x)\right\rvert^{2}f_{\infty}(x;\theta_{0})\,\mathrm{d}x=0\right\}.$ (3.36)
+Finally, by the identifiability condition in Assumption˜3.1(iii) we get that $\Omega=\{\theta_{0}\}$, and the desired result follows. ∎
 3.1.2 Numerical examples
-Figure 1: Comparison between the MLE $\widehat{\theta}_{T}$ and the linearized MLE $\widetilde{\theta}_{T}$ for different values of the final time of observation $T\in[0,10^{3}]$, considering the quadratic interaction potential and two different examples of confining potentials.
+Figure 1: Comparison between the MLE $\widehat{\theta}_{T}$ and the linearized MLE $\widetilde{\theta}_{T}$ for different values of the final time of observation $T\in[0,10^{3}]$, considering the quadratic interaction potential and two different examples of confining potentials, when the unknown coefficient is in the confining potential.
+Figure 2: Comparison between the MLE $\widehat{\theta}_{T}$ and the linearized MLE $\widetilde{\theta}_{T}$ for different values of the final time of observation $T\in[0,10^{3}]$, considering the quadratic interaction potential and two different examples of confining potentials, when the unknown coefficient is in the interaction potential.
 We now implement the linearized MLE and test it on two numerical examples.
@@
 $\widetilde{\theta}_{T}=-\frac{\int_{0}^{T}X_{t}\,\mathrm{d}X_{t}}{\int_{0}^{T}X_{t}^{2}\,\mathrm{d}t}-1.$ (3.43)
-In the first plot in Figure 1, we compare the two estimators for different values of the final time of observation $T\in[0,10^{3}]$, and notice that, except for a short time interval, the difference between $\widehat{\theta}_{T}$ and $\widetilde{\theta}_{T}$ is negligible.
+In the first plot in Figure˜1, we compare the two estimators for different values of the final time of observation $T\in[0,10^{3}]$, and notice that, except for a short time interval, the difference between $\widehat{\theta}_{T}$ and $\widetilde{\theta}_{T}$ is negligible.
 In particular, they both provide an accurate approximation of the exact unknown.
@@
 $\mathrm{d}X_{t}$ $=-\theta(X_{t}^{3}-X_{t})\,\mathrm{d}t-(X_{t}-\operatorname{\mathbb{E}}[X_{t}])\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},$ (3.45)
-$\mathrm{d}Y_{t}$ $=-\theta X_{t}^{3}\,\mathrm{d}t-(1-\theta)X_{t}\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t}.$
+$\mathrm{d}Y_{t}$ $=-\theta Y_{t}^{3}\,\mathrm{d}t-(1-\theta)Y_{t}\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t}.$
 Assuming to know, in addition to the trajectory $(X_{t})_{t\in[0,T]}$, also its expectation at all times $(\operatorname{\mathbb{E}}[X_{t}])_{t\in[0,T]}$, we can write the MLE as
@@
 On the contrary, for the linearized MLE, only the trajectory of a single particle is required.
-The behavior of the two estimators for different values of the final time of observation $T\in[0,10^{3}]$ is shown in the second plot in Figure 1.
+The behavior of the two estimators for different values of the final time of observation $T\in[0,10^{3}]$ is shown in the second plot in Figure˜1.
 As expected, the two estimators provide us with almost identical values for the estimated parameter, after a short initial transient time.
+We finally remark that similar results are obtained if the unknown parameter to be estimated appears in the interaction potential.
+Specifically, consider the mean field Ornstein–Uhlenbeck process together with its linearized process
+$\mathrm{d}X_{t}$ $=-X_{t}\,\mathrm{d}t-\theta(X_{t}-\operatorname{\mathbb{E}}[X_{t}])\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},$ (3.48)
+$\mathrm{d}Y_{t}$ $=-(\theta+1)Y_{t}\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t}.$
+In this case we have $\operatorname{\mathbb{E}}[X_{t}]=\operatorname{\mathbb{E}}[X_{0}]e^{-t}$ and the MLE estimators are
+$\widehat{\theta}_{T}$ $=-\frac{\int_{0}^{T}(X_{t}-\operatorname{\mathbb{E}}[X_{0}]e^{-t})\,\mathrm{d}X_{t}+\int_{0}^{T}X_{t}(X_{t}-\operatorname{\mathbb{E}}[X_{0}]e^{-t})\,\mathrm{d}t}{\int_{0}^{T}(X_{t}-\operatorname{\mathbb{E}}[X_{0}]e^{-t})^{2}\,\mathrm{d}t},$ (3.49)
+$\widetilde{\theta}_{T}$ $=-\frac{\int_{0}^{T}X_{t}\,\mathrm{d}X_{t}}{\int_{0}^{T}X_{t}^{2}\,\mathrm{d}t}-1.$
+On the other hand, regarding the bistable confining potential example, the linear and nonlinear processes satisfy
+$\mathrm{d}X_{t}$ $=-(X_{t}^{3}-X_{t})\,\mathrm{d}t-\theta(X_{t}-\operatorname{\mathbb{E}}[X_{t}])\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},$ (3.50)
+$\mathrm{d}Y_{t}$ $=-Y_{t}^{3}\,\mathrm{d}t-(\theta-1)Y_{t}\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t},$
+and the corresponding MLE estimators read
+$\widehat{\theta}_{T}$ $=-\frac{\int_{0}^{T}(X_{t}-\operatorname{\mathbb{E}}[X_{t}])\,\mathrm{d}X_{t}+\int_{0}^{T}(X_{t}-\operatorname{\mathbb{E}}[X_{t}])(X_{t}^{3}-X_{t})\,\mathrm{d}t}{\int_{0}^{T}(X_{t}-\operatorname{\mathbb{E}}[X_{t}])^{2}\,\mathrm{d}t},$ (3.51)
+$\widetilde{\theta}_{T}$ $=-\frac{\int_{0}^{T}X_{t}\,\mathrm{d}X_{t}+\int_{0}^{T}X_{t}(X_{t}^{3}-X_{t})\,\mathrm{d}t}{\int_{0}^{T}X_{t}^{2}\,\mathrm{d}t}.$
+The numerical results are reported in Figure˜2, where we observe that the linear and nonlinear MLE estimators agree and tend to coincide when the final time of observation increases.
 3.2 The diffusive-mean field limit
@@
 As discussed in that paper, the two limits of sending $\varepsilon\to 0$ and $N\to\infty$ commute only when the mean field dynamics on the torus does not exhibit phase transitions.
-Here, we study only the “easy” order of taking the two limits, first $N\to\infty$ and then sending $\varepsilon$ to 00, and only under the assumption that the mean field SDE has a unique steady state.
+Here, we study only the “easy” order of taking the two limits, first $N\to\infty$ and then sending $\varepsilon$ to 0, and only under the assumption that the mean field SDE has a unique steady state.
 A full analysis of the problem based on the linearized SDE, for both the overdamped and the underdamped dynamics, will be presented elsewhere.
@@
 Theorem 3.10.
-Let $Y_{t}$ be the solution of the SDE (3.58), and let $\widetilde{Y}^{\varepsilon}_{t}=\varepsilon Y^{\varepsilon}_{t/\varepsilon^{2}}$. Then, it holds
+Let $Y_{t}$ be the solution of the SDE (3.62), and let $\widetilde{Y}^{\varepsilon}_{t}=\varepsilon Y_{t/\varepsilon^{2}}$. Then, it holds
 $\lim_{\varepsilon\to 0}\widetilde{Y}^{\varepsilon}_{t}=\sqrt{2\beta^{-1}\mathcal{D}}B_{t},\qquad\text{in law},$ (3.70)
@@
 Proof.
-Reasoning as in the proof of Theorem 3.9, we obtain
+Reasoning as in the proof of Theorem˜3.9, we obtain
 $\widetilde{Y}^{\varepsilon}_{t}=\varepsilon Y_{t/\varepsilon^{2}}=\varepsilon\left(Y_{0}+\Phi(Y_{0})-\Phi(Y_{t/\varepsilon^{2}})\right)+\varepsilon\sqrt{2\beta^{-1}}\int_{0}^{t/\varepsilon^{2}}(I+\nabla\Phi(Y_{s}))\,\mathrm{d}B_{s},$ (3.71)
@@
 3.2.2 Numerical example
-Figure 2: Numerical verification of the CLTs in equation (3.57) and Theorem 3.9, considering sinusoidal confining and interaction potentials.
-We numerically verify in a one-dimensional test case the CLTs for the nonlinear and linear processes $X_{t}$ and $Y_{t}$ given in equations (3.57) and Theorem 3.9, respectively.
+Figure 3: Numerical verification of the CLTs in equation (3.61) and Theorem˜3.9, considering sinusoidal confining and interaction potentials.
+We numerically verify in a one-dimensional test case the CLTs for the nonlinear and linear processes $X_{t}$ and $Y_{t}$ given in equations (3.61) and Theorem˜3.9, respectively.
 We consider the setting of [22, Lemma 1.26] below the phase transition, i.e., we set $W(x)=-\cos(2\pi x)$, $V(x)=-\xi\cos(2\pi x)$ with $\xi=1/2$, and $\beta=1$.
@@
 $\,\mathrm{d}Y_{t}=-2\pi\beta^{-1}A\sin(2\pi Y_{t})\,\mathrm{d}t+\sqrt{2\beta^{-1}}\,\mathrm{d}B_{t}.$ (3.80)
-In Figure 2 we show the histograms of $X_{t}/\sqrt{t}$ and $Y_{t}/\sqrt{t}$ with $t=10^{3}$ for $500$ different realizations of the Brownian motion.
+In Figure˜3 we show the histograms of $X_{t}/\sqrt{t}$ and $Y_{t}/\sqrt{t}$ with $t=10^{3}$ for 500500500 different realizations of the Brownian motion.
 We remark that, differently from equation (3.80) that can be simulated, the McKean SDE for $X_{t}$ must be approximated by the interacting particle system (1.1), for which we use $N=250$ particles.